### v4

# From Unaligned to Aligned: Scaling Multilingual LLMs with Multi-Way Parallel Corpora

Abstract
Continued pretraining and instruction tuning on large-scale multilingual data have proven to be effective in scaling large language models (LLMs) to low-resource languages. However, the unaligned nature of such data limits its ability to effectively capture cross-lingual semantics. In contrast, multi-way parallel data, where identical content is aligned across multiple languages, provides stronger cross-lingual consistency and offers greater potential for improving multilingual performance. In this paper, we introduce a large-scale, high-quality multi-way parallel corpus, TED2025, based on TED Talks. The corpus spans 113 languages, with up to 50 languages aligned in parallel, ensuring extensive multilingual coverage. Using this dataset, we investigate best practices for leveraging multi-way parallel data to enhance LLMs, including strategies for continued pretraining, instruction tuning, and the analysis of key influencing factors. Experiments on six multilingual benchmarks show that models trained on multi-way parallel data consistently outperform those trained on unaligned multilingual data. [^1] https://github.com/yl-shen/multi-way-llm
From Unaligned to Aligned: Scaling Multilingual LLMs with Multi-Way Parallel Corpora
Yingli Shen1thanks: Equal contribution., Wen Lai2,3∗, Shuo Wang1, Ge Gao4 Kangyang Luo1, Alexander Fraser2,3, Maosong Sun1,5thanks: Corresponding author. 1 Department of Computer Science and Technology, Tsinghua University 2 Technical University of Munich  3 Munich Center for Machine Learning 4 Minzu University of China  5 Institute for AI, Tsinghua University syl@mail.tsinghua.edu.cn, wen.lai@tum.de
1 Introduction
Large language models (LLMs) have demonstrated remarkable performance on various tasks in high-resource languages (Huang et al., 2024; Qin et al., 2024). However, their performance still lags behind in low-resource languages (Huang et al., 2023; Lai et al., 2023a). To bridge this gap, recent efforts have focused on continued pretraining (Ji et al., 2024; Groeneveld et al., 2024) and instruction tuning (Lai et al., 2024; Üstün et al., 2024), utilizing large-scale unaligned multilingual text. However, these methods do not take full advantage of the explicit many-to-many alignments present in multi-way parallel corpora, which have been shown to improve cross-lingual representations in multilingual NLP (Qi et al., 2018; Freitag and Firat, 2020; Xu et al., 2022; Wu et al., 2024). More recently, Mu et al. (2024) demonstrated that multi-way parallel inputs can also enhance in-context learning (Dong et al., 2024). However, scaling multilingual LLMs using multi-way parallel data remains underexplored.
Dataset Source #Langs #Domains Max Parallelism Type Collection Method Date Range Open-source?
Bible Christodouloupoulos and Steedman (2015) Bible 100 1 55 Training Human Translator 2015
UN Corpus Ziemski et al. (2016) UN 6 1 6 Training Human Translator 2016
MMCR4NLP Dabre and Kurohashi (2017) Mix 59 5 13 Training Mix Collection 2017
GCP Corpus Imamura and Sumita (2018) Speech 10 4 10 Training Machine Translation 2018
FLORES-101 Goyal et al. (2022) Wiki 101 10 101 Evaluation Human Translator 2022
FLORES-200 Costa-Jussà et al. (2022) Wiki 204 10 204 Evaluation Human Translator 2022
BPCC Gala et al. (2023) Many 22 13 22 Training Mix Collection 2023
XDailyDialog Liu et al. (2023b) DailyDialog 4 / 4 Training Human Translator 2023
MWccMatrix Thompson et al. (2024) Common Crawl 90 / / Training Crawl 2024
TED2018 Qi et al. (2018) TED 58 / 58 Training Human Translator 2018
TED2020 Reimers and Gurevych (2020) TED 108 / / Training Human Translator 2020
Ours (TED2025) TED 113 352 50 Training Human Translator 2025
Table 1: Comparison of existing multi-way parallel corpora and our constructed TED2025, highlighting key attributes such as the data source, the number of languages (#Langs), the number of domains (#Domains), the maximum parallelism supported, the type of data (training or evaluation), the collection method, and whether the corpus is open-source.
Existing multi-way parallel datasets typically cover only a limited number of languages, domains and levels of parallelism (see Table 1). In contrast, TED Translators [^2] https://www.ted.com/participate/translate, a global community translating TED talk transcripts into over 100 languages, provides consistently high-quality, human-verified translations and serves as an ideal source for a large-scale multi-way parallel corpus. However, the largest TED-based datasets (Qi et al., 2018; Reimers and Gurevych, 2020) have not been updated since 2020, limiting their utility for LLM training and potentially exacerbating hallucinations (Ji et al., 2023).
To address these limitations, we introduce TED2025, a large-scale, high-quality multi-way parallel corpus derived from the latest TED talks. TED2025 covers 113 languages, 352 domain labels, and supports up to 50-way parallelism. Compared to existing resources, it offers more frequent updates and significantly broader coverage across both languages and domains, thereby strengthening the data foundation for multilingual LLM training.
Utilizing TED2025, we investigate three key research questions:
RQ1: How does fine-tuning on multi-way parallel data compare to training on unaligned multilingual text in terms of zero-shot cross-lingual transfer and representation alignment?
RQ2: Which strategies for selecting parallelism in multi-way parallel data (e.g., degree of parallelism and language subsets) lead to the greatest improvements in multilingual LLM performance?
RQ3: Which instruction-tuning objectives can most effectively leverage the advantages of multi-way parallel data?
We perform a comprehensive evaluation across six multilingual benchmarks to assess the benefits of using multi-way parallel data for scaling multilingual LLMs. Our results reveal that, at an equal data scale, fine-tuning on multi-way parallel data consistently outperforms training on unaligned multilingual text for both low-resource and high-resource languages (Section 4). Additionally, we identify the most effective configurations of parallelism (Section 5). Furthermore, we investigate how different instruction-tuning objectives impact LLM performance and cross-domain robustness (Section 6).
In summary, our contributions are as follows: (i) We construct TED2025, a 50-way parallel corpus derived from recent TED talk transcripts, covering 113 languages and 352 domains. (ii) To the best of our knowledge, this is the first work to leverage multi-way parallel data for scaling multilingual LLMs. We present a systematic comparison of multilingual LLM fine-tuning using multi-way versus unaligned data, analyzing their effects on zero-shot transfer and cross-lingual representation alignment. (iii) We explore instruction-tuning objectives specifically designed for multi-way parallel data and provide practical recommendations for optimizing multilingual LLM performance.
2 Related Work
Multi-Way Parallel Corpora.
Datasets containing the same content across multiple languages (typically more than two) are known as multi-way parallel corpora. These corpora have demonstrated substantial benefits for machine translation (Freitag and Firat, 2020; Xu et al., 2022; Wu et al., 2024) and cross-lingual representation alignment (Tran et al., 2020). Existing methods for constructing such corpora include mining comparable texts (Resnik et al., 1999), aligning independently collected monolingual corpora via translation pivots (Thompson et al., 2024), extracting multi-way subsets from large bilingual collections (Ramesh et al., 2022), and harvesting multilingual web crawls (Resnik and Smith, 2003; Qi et al., 2018). However, many of these resources are limited in terms of language and domain coverage. In contrast, we construct a multi-way parallel corpus derived from recent TED Talk transcripts, offering a broader and more diverse set of languages and domains.
Scaling Multilingual LLMs.
The multilingual capabilities of LLMs have been significantly enhanced through two complementary strategies: continued pretraining on diverse multilingual corpora and multilingual instruction tuning. Continued pretraining on unaligned multilingual data has improved both in-language fluency and cross-lingual transfer (Ji et al., 2024; Groeneveld et al., 2024), while instruction tuning with human-curated multilingual prompts has boosted task performance across a wide range of languages (Lai et al., 2024; Üstün et al., 2024). More recently, Mu et al. (2024) demonstrated that incorporating multi-way parallel examples into in-context prompts leads to further gains in zero-shot transfer. Building on these insights, we systematically fine-tune multilingual LLMs on large-scale multi-way parallel data and quantify their impact compared to conventional unaligned approaches.
3 Experimental Setup
Figure 1: Distribution of sentence counts (line chart, left y-axis, log scale) and parallelism spans (bar chart, right y-axis, ratio) across languages (x-axis) in the TED2025 corpus. The parallelism spans, with a notable concentration between 21 and 30 languages, and high range even for low-resource languages.
TED2025.
We introduce TED2025, a new multi-way parallel corpus derived from the latest TED Talk transcripts. It encompasses 113 languages with up to 50-way parallelism, making it one of the largest and most diverse resources for multilingual fine-tuning. Figure 1 illustrates the total number of sentences and the distribution of parallelism spans across languages in TED2025. Figure 2 compares translation quality, as measured by COMET-QE (Rei et al., 2020), across TED2025 (which includes 4,765 language pairs [^3] For language pairs with more than 10,000 sentence pairs, we randomly sample 10,000.) and other existing multi-way datasets: TED2018 (Qi et al., 2018), TED2020 (Reimers and Gurevych, 2020), and MWccMatrix (Thompson et al., 2024). Additional dataset statistics and construction details are provided in Appendix A.
We observe that: (1) the most common parallelism span in TED2025 ranges from 21 to 30 languages. Notably, many low-resource languages also achieve high degrees of parallelism, providing a solid foundation for multilingual research. (2) TED2025 contains significantly more high-quality translations (with a COMET-QE score greater than 60) compared to previous multi-way corpora. Unlike TED2020, which segments English based on punctuation, or MWccMatrix, which relies on LASER score (Artetxe and Schwenk, 2019), we use human-provided timestamps to generate cleaner and more reliable sentence alignments.
Figure 2: Comparison of translation quality between TED2025 and existing multi-way datasets, including TED2018 Qi et al. (2018), TED2020 Reimers and Gurevych (2020), MWccMatrix Thompson et al. (2024), using COMET-QE score.
Training.
To isolate the effects of multi-way parallel data, we conduct both continued pretraining (Parmar et al., 2024) and instruction tuning (Zhang et al., 2023) on TED2025. We experiment with two model families and sizes: LLaMA-3.1-8B / LLaMA-3.1-8B-Instruct and Qwen-2.5-14B / Qwen-2.5-14B-Instruct (available on Hugging Face [^4] https://huggingface.co). To make fine-tuning feasible, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) instead of performing full parameter updates. Full hyperparameter settings and training configurations are provided in Appendix B.
Benchmark Task #Langs Metric
MMMLU Understanding 14 Acc
XCOPA Reasoning 11 Acc
FLORES-101 Generation 101 BLEU/COMET
FLORES-200 Generation 204 BLEU/COMET
xIFEval Instruction Following 17 Acc
SIB-200 Text Classification 204 Acc
Table 2: Overview of evaluation benchmarks, including task types, the number of languages (#Langs) involved, and the metrics used for assessment.
Evaluation and Metrics.
We evaluate our models on five widely adopted multilingual benchmarks in a zero-shot setting, covering a range of tasks: understanding (MMMLU), reasoning (XCOPA; Ponti et al., 2020), generation (FLORES-101; Goyal et al., 2022 and FLORES-200; Costa-Jussà et al., 2022), instruction following (xIFEval; Huang et al., 2025), and text classification (SIB-200; Adelani et al., 2024). Table 2 summarizes these benchmarks along with their associated evaluation metrics. Additionally, we categorize the languages in each benchmark as low-resource or high-resource based on the classification in Costa-Jussà et al. (2022).
4 Effectiveness of Multi-Way Corpora
We investigate the impact of training on multi-way parallel data on the performance of a multilingual LLM across three key dimensions: downstream performance on multilingual benchmarks (Section 4.1), zero-shot cross-lingual transfer to unseen languages (Section 4.2) and cross-lingual alignment of representations within the model’s internal embeddings (Section 4.3).
For fairness, we fix the total continued pretraining data at 5 million tokens (5M) and evaluate two LLM backbones (LLaMA-3-8B and Qwen-2.5-14B) under three conditions: (1) Multi-Way: Pretraining on our multi-way parallel corpus (TED2025); (2) Unaligned: Pretraining on an unaligned multilingual corpus [^5] Unaligned data is sourced from DCAD-2000 Shen et al. (2025) rather than TED2025, due to considerations related to scalability and domain coverage. (DCAD-2000; Shen et al., 2025); (3) Baseline: The original pretrained checkpoint without additional data.
MMMLU XCOPA FLORES-101 (Eng-X) FLORES-101 (X-Eng) xIFEval
BLEU COMET BLEU COMET
low high low high low high low high low high low high low high
Baseline 18.27 33.72 23.46 34.29 6.03 11.67 57.15 61.03 13.37 22.49 75.24 82.32 17.14 24.43
Unaligned 19.64 36.26 24.62 34.76 6.12 11.78 57.51 62.11 13.84 22.74 75.82 82.58 17.28 24.44
Multi-Way 22.48 41.38 27.58 57.22 6.32 12.08 58.06 67.44 14.45 25.03 76.25 86.43 18.79 27.41
(a) LLaMA-3-8B
MMMLU XCOPA FLORES-101 (Eng-X) FLORES-101 (X-Eng) xIFEval
BLEU COMET BLEU COMET
low high low high low high low high low high low high low high
Baseline 35.24 49.55 62.25 72.00 7.45 11.05 57.22 67.16 16.54 20.24 67.23 74.29 27.63 32.40
Unaligned 35.61 51.32 62.59 74.06 7.62 11.60 57.85 70.85 16.86 21.02 67.61 75.97 27.92 35.54
Multi-Way 36.64 55.81 63.24 79.52 8.07 13.11 58.94 80.56 17.36 23.26 68.59 81.33 28.64 40.95
(b) Qwen-2.5-14B
Table 3: Performance (%) comparison of different models across multilingual benchmarks. The Multi-Way approach with blue background demonstrates consistent superiority in both low-resource (left columns) and high-resource (right columns) scenarios.
4.1 Downstream Performance
We evaluate all variants in a zero-shot setting across four benchmarks—MMMLU, XCOPA, FLORES-101, and xIFEval—that cover understanding, reasoning, generation, and instruction following. The full results are reported in Table 3. Our findings show that, across all tasks, Multi-Way consistently outperforms both Baseline and Unaligned across both low- and high-resource languages. For example, on MMMLU, Multi-Way achieves accuracies of 22.48/41.38 in low/high-resource languages, compared to 18.27/33.72 for the Baseline and 19.64/36.26 for Unaligned. Similar improvements are observed on FLORES-101 and xIFEval. These results demonstrate that multi-way alignment provides stronger cross-lingual supervision, thereby enhancing both discriminative and generative capabilities. Although our main experiments focus on English–X translation directions, we also evaluate the trained models on non-English-centric translation pairs. The results, which show that models trained on aligned data significantly outperform those trained on unaligned data even for non-English-centric tasks, are provided in Appendix C.
Figure 3: Cross-lingual transfer performance comparison between Baseline, Unaligned and Multi-Way pretraining on the FLORES-200 benchmark with BLEU (bar chart, left y-axis) and COMET (line chart, right y-axis) for LLaMA-3-8B and Qwen-2.5-14B models.
4.2 Zero-Shot Cross-Lingual Transfer
[…]
LLaMA-3-8B Qwen-2.5-14B
Baseline Unaligned Multi-Way Baseline Unaligned Multi-Way
Cosine ($\uparrow$) 0.27 0.27-0.00 0.30+0.03 0.29 0.27-0.02 0.32+0.03
CKA ($\uparrow$) 0.54 0.54-0.00 0.60+0.06 0.56 0.57+0.01 0.63+0.07
Retrieval ($\uparrow$) P@1 0.09 0.07-0.02 0.13+0.04 0.12 0.14+0.02 0.19+0.07
P@5 0.24 0.23-0.01 0.27+0.03 0.26 0.28+0.02 0.33+0.07
P@10 0.35 0.33-0.02 0.39+0.04 0.36 0.38+0.02 0.42+0.06
SVCCA ($\uparrow$) 0.55 0.55-0.00 0.61+0.06 0.57 0.58+0.01 0.63+0.06
Table 4: Cross-lingual representation alignment results. Improvement margins (colored red/blue) are shown relative to Baseline. Bolded Multi-Way results with red annotations indicate consistent improvements, particularly in cross-lingual retrieval tasks (e.g., +0.07 P@1 improvement for Qwen-2.5-14B).
4.3 Cross-Lingual Representation Alignment
We further analyze the alignment of internal representations across models. To be specific, for 32 randomly selected aligned languages (with 100 sentences each) from TED2025, we compute the following four metrics: average cosine similarity between parallel sentence embeddings, Centered Kernel Alignment (CKA) between representation matrices (Kornblith et al., 2019), Cross-lingual sentence retrieval accuracy at P@1, P@5, and P@10 (Conneau et al., 2017) and SVCCA score (Raghu et al., 2017).
Table 4 shows that Multi‑Way outperforms the other models, yielding higher CKA and better retrieval accuracy. Figure 4 further corroborates these results: Multi‑Way demonstrates denser SVCCA alignments, particularly for linguistically distant language pairs. These metrics confirm that multi-way pretraining promotes a more coherent, language-agnostic embedding space, which drives the observed improvements in downstream performance and cross-lingual transfer.
5 Impact Factors
In Section 4, we demonstrate that using multi-way parallel data can significantly enhance the multilingual capabilities of LLMs. To better understand the factors driving this improvement, we analyze two key aspects, while keeping the total pretraining fixed at 5 million tokens: (1) Degree of parallelism: the number of languages aligned in each training example (Section 5.1). (2) English as a pivot: the impact of including versus excluding English in multi-way groups (Section 5.2). (3) Language combinations: the role of language family composition in training data (Section 5.3). (4) Training data size: how the amount of pretraining data affects model performance (Section 5.4).
Figure 4: SVCCA alignment comparison between the Multi-Way, Unaligned and Baseline models across 32-way language pairs.
5.1 Degree of Parallelism
[…]
Figure 5: Performance (%) of continued pretraining models on downstream tasks with varying degrees of parallelism.
5.2 English as Pivot
English is widely used as a pivot language in multilingual MT and NLP Kim et al. (2019); Mallinson et al. (2018); Lai et al. (2023b). To explore its impact, we form five groups [^6] We provide the configurations of each group in Table 10 of Appendix D, each with both English-included and English-excluded variants across ten languages. Table 5 compares their performance across various tasks.
In tasks involving understanding and reasoning, groups that include English consistently outperform those in other language groups by an average of 2–4 percentage points. This suggests that English, as a high-resource “semantic anchor”, helps stabilize embedding alignment and facilitates transfer learning, particularly on complex tasks. Interestingly, for machine translation (FLORES-101) and some instruction-following tasks (xIFEval), English-inclusive groups show slightly lower performance. We attribute this to two primary factors: (1) English occupies tokens that could otherwise be used to align non-English language pairs directly. (2) The model may overly rely on English as an intermediary, which reduces its ability to directly transfer knowledge between non-English languages. These findings highlight that English’s role as a pivot language is task-dependent. While it can enhance semantic coherence in understanding and reasoning tasks, it may hinder direct multilingual transfer in generative tasks. Consequently, the inclusion of English in multilingual pretraining should be carefully considered based on the target application.
MMMLU XCOPA
FLORES
(Eng-X)
FLORES
(X-Eng)
xIFEval
Group 1 with 23.17 30.93 5.80 13.75 23.39
w/o 20.84 30.07 6.63 14.56 24.14
Group 2 with 22.19 35.33 6.74 13.88 22.19
w/o 18.51 31.60 7.13 14.18 22.19
Group 3 with 26.47 36.73 6.40 13.48 22.69
w/o 23.83 35.78 6.38 14.00 22.91
Group 4 with 23.42 33.84 6.20 12.11 23.66
w/o 20.26 30.84 6.67 14.76 23.67
Group 5 with 23.89 39.64 6.96 13.07 23.19
w/o 20.39 34.15 7.83 14.79 23.85
Table 5: Performance (%) comparison of models with and without (w/o) English across five different language groupings.
MMMLU XCOPA FLORES-101 (Eng-X) FLORES-101 (X-Eng) xIFEval
BLEU COMET BLEU COMET
low high low high low high low high low high low high low high
Baseline 41.96 46.68 64.59 66.79 5.51 10.50 56.95 60.02 14.98 18.41 68.29 73.51 35.99 38.56
MT 45.28 51.06 68.17 69.38 11.26 13.25 63.03 65.61 22.25 21.60 70.76 75.41 41.72 44.52
CLTS 39.99 45.44 62.87 66.47 3.63 9.90 56.48 59.28 13.25 16.60 66.89 73.15 34.91 38.38
MTC 40.68 44.49 64.19 65.16 5.02 9.08 56.04 57.84 14.38 18.16 67.86 73.12 34.03 38.01
CLP 42.49 47.01 65.41 68.42 6.38 11.46 57.23 61.28 15.76 19.27 69.87 73.91 36.17 40.07
MT + CLTS 42.23 47.94 65.00 68.77 6.19 10.51 58.53 60.33 16.83 18.98 68.70 74.91 36.47 40.30
MT + MTC 42.69 47.54 65.12 66.83 6.35 10.73 57.53 60.24 15.13 18.45 69.28 74.00 36.10 39.03
MT + CLP 43.20 49.07 67.29 68.47 7.31 10.51 58.40 62.64 17.75 20.77 69.18 74.47 38.49 39.67
CLTS + MTC 41.78 45.17 63.21 65.62 5.38 9.77 55.19 58.21 14.67 16.59 67.45 72.35 34.57 36.63
CLTS + CLP 42.82 46.74 65.30 67.12 5.58 10.88 57.84 60.41 15.41 19.33 68.84 73.58 36.99 38.83
MTC + CLP 42.62 46.70 65.16 67.56 6.48 10.84 57.74 60.98 15.29 19.22 68.76 73.77 36.87 38.96
MT + CLTS + MTC 41.03 46.59 64.15 66.05 5.14 10.35 56.14 59.71 14.64 18.06 67.64 73.39 35.24 38.14
MT + CLTS + CLP 42.72 47.67 64.83 67.57 6.17 10.94 57.78 60.11 15.96 19.16 68.67 74.03 36.54 38.77
MT + MTC + CLP 42.33 46.72 65.26 67.58 5.92 10.98 57.93 60.76 15.38 18.81 68.98 73.55 36.22 39.15
CLTS + MTC + CLP 41.56 46.65 64.16 66.45 4.67 10.38 56.20 59.63 14.95 17.62 68.00 73.03 35.27 38.21
MT + CLTS + MTC + CLP 43.07 47.67 66.64 67.38 7.59 12.42 57.75 60.79 17.62 19.62 71.13 75.24 36.95 40.52
(a) LLaMA-3-8B-Instruct
MMMLU XCOPA FLORES-101 (Eng-X) FLORES-101 (X-Eng) xIFEval
BLEU COMET BLEU COMET
low high low high low high low high low high low high low high
Baseline 63.61 68.14 78.76 82.22 7.92 12.79 61.49 66.64 16.76 20.48 68.16 70.32 47.68 52.40
MT 68.52 72.83 82.95 86.95 11.90 16.60 64.96 70.90 20.43 26.22 72.25 73.51 53.24 55.83
CLTS 61.28 67.61 76.77 79.74 7.20 10.70 60.90 64.50 15.75 18.21 67.01 69.83 45.73 51.83
MTC 62.59 65.63 76.87 81.46 6.07 12.29 61.48 65.61 13.84 19.71 67.12 68.65 47.53 50.53
CLP 64.28 70.06 79.69 83.02 9.54 13.10 62.18 68.35 17.54 21.68 69.24 70.68 48.61 52.66
MT + CLTS 63.93 69.53 78.94 82.75 8.73 13.91 61.87 67.09 18.20 21.50 68.45 71.90 48.16 54.14
MT + MTC 65.33 69.12 80.27 83.34 8.64 13.94 61.71 67.42 17.87 22.27 68.60 72.01 49.46 53.53
MT + CLP 64.95 70.51 80.68 83.52 10.30 13.65 62.37 69.32 18.43 22.19 70.01 70.80 48.85 53.15
CLTS + MTC 62.67 67.19 78.44 81.33 7.61 12.60 61.04 65.90 16.25 19.93 67.47 70.00 46.85 51.63
CLTS + CLP 64.38 68.65 78.89 82.40 8.65 13.57 61.75 67.57 17.73 20.81 68.65 71.30 47.82 53.04
MTC + CLP 64.04 68.92 79.16 82.56 8.81 13.05 62.26 67.34 16.97 21.31 68.98 71.20 48.06 53.32
MT + CLTS + MTC 64.13 68.55 79.51 82.63 8.26 13.30 62.24 67.01 17.69 21.29 68.38 70.74 48.54 52.80
MT + CLTS + CLP 64.50 68.82 80.43 83.22 9.05 13.93 62.38 67.56 18.66 22.01 68.76 71.62 48.96 53.74
MT + MTC + CLP 64.71 69.52 80.26 83.32 9.15 13.37 62.82 67.66 17.75 21.70 69.33 71.50 48.93 53.41
CLTS + MTC + CLP 63.81 69.00 79.70 82.52 8.69 13.48 62.25 67.04 17.55 20.75 68.23 70.96 47.68 52.55
MT + CLTS + MTC + CLP 64.94 68.19 80.25 82.42 9.87 13.56 62.86 67.04 16.94 21.69 68.17 71.34 49.03 53.72
(b) Qwen-2.5-14B-Instruct
Table 6: Downstream task performance (%) of LLaMA-3-8B-Instruct and Qwen-2.5-14B-Instruct models trained with different instruction objectives, including MT, CLTS, MTC and CLP.
5.3 Language Combinations
In Section 5.2, we investigate the influence of including English in sampling combinations on model performance. In this section, we extend that analysis by investigating an alternative sampling strategy: whether the selected languages belong to the same language family. To this end, we construct four language groups. Groups 1 and 2 consist of languages from the same language family, whereas Groups 3 and 4 include languages from diverse families. The specific language configurations for each group are detailed in Appendix D (Table 10).
Figure 6 illustrates the impact of language family composition on model performance. The results indicate that sampling from cross-family language combinations more effectively leverages the benefits of multi-way parallel corpora, resulting in greater improvements in multilingual task performance. In contrast, sampling languages from the same family yields only marginal gains. This may be attributed to the structural and lexical similarities among related languages, which can introduce redundancy during training and limit the model’s ability to generalize across typologically diverse languages.
Figure 6: Impact of language family composition on model performance.
Figure 7: Impact of training data size on model performance across different token amounts (ranging from 10K to 1B) sampled from the TED2025 dataset.
5.4 Training Data Size
Figure 7 presents the impact of training data size on model performance. We conduct experiments by randomly sampling varying amounts of tokens—10K, 50K, 100K, 500K, 1M, 5M, 10M, 50M, 100M, 500M, and 1B—from the constructed TED2025 dataset. The results demonstrate that model performance is notably constrained when trained on smaller datasets (typically under 100K tokens). However, as the size of training data increases, performance consistently improves across all evaluated tasks.
For MMMLU and XCOPA, performance exhibits early gains with additional data but plateaus beyond a certain threshold. This trend likely reflects the nature of these tasks, which emphasize general language understanding and reasoning. Once the model acquires the necessary core linguistic and world knowledge, the marginal gains from further data diminish. Interestingly, performance on FLORES and xIFEval continues to improve with increasing data volume. These tasks, which involve cross-lingual understanding and translation—particularly for low-resource languages or semantically nuanced alignments—appear to benefit more substantially from large-scale data. This suggests that extensive training data is crucial for enhancing translation quality and evaluation accuracy in such settings.
6 Instruction Tuning
Task Prompt
[…]
Similarity: {sim_score}.
[…]
Categories: {target_domain}.
[…]
{tgt_lang} Sentence: {tgt_txt}.
Table 7: Instruction prompts used for four multilingual tasks: machine translation (MT), cross-lingual text similarity (CLTS), multilingual text classification (MTC), and cross-lingual paraphrasing (CLP). Each prompt is designed to reflect the task’s specific objective and structure.
In this section, we further investigate whether instruction tuning can also enhance multilingual performance effectively. Specifically, we address the following key questions [^7] Additionally, we explore the cross-lingual alignment of representations in instruction tuned models (Appendix E.): (1) Which of the different instruction fine-tuning objectives, built on multi-way parallel data, is most effective? (2) Do models trained with multi-way parallel data exhibit better generalization across domains?
6.1 Instruction Tuning Objectives
Using our constructed multi-way parallel data (TED2025), we define four instruction tasks: machine translation (MT), cross-lingual text similarity (CLTS), multilingual text classification (MTC), and cross-language paraphrasing (CLP). Table 6 reports their impact on downstream benchmarks. Table 7 summarizes the prompt templates and output formats for each task.
We observe that the improvements in MT are the largest and most stable in both high-resource and low-resource languages. This can primarily be attributed to the fact that, as a token-level supervised generation task, translation strengthens cross-lingual syntactic and semantic consistency, making it a particularly robust task for broad multilingual generalization. In contrast, CLTS and MTC show smaller drops across different tasks, which we attribute to the coarser granularity of similarity judgments that may not provide fine-grained alignment signals. Moreover, discrete class labels lack the expressive power to capture subtle semantic distinctions. Although CLP shows similar advantages to MT in generation tasks, its advantages are narrower in scope and cannot be widely generalized.
Interestingly, the combination of tasks did not significantly improve performance, which may be due to several reasons: First, the objectives of MT and CLP differ; while MT emphasizes accuracy, CLP focuses more on the diversity of expression, which may make it difficult for the model to balance these two goals during training, thereby affecting performance. Second, interference between multiple tasks may arise, making it challenging for the model to focus on optimizing the optimal goal of each task, particularly when the task objectives are too similar, amplifying the interference effect. Furthermore, there is overlap in task design between MT and CLP, which may cause the model to encounter redundant training signals when processing both tasks, preventing it from fully utilizing the unique value of each task. Finally, joint training of multiple tasks increases the complexity of model training, potentially leading to unstable gradient updates and affecting the model’s convergence.
6.2 Cross-Domain Generalization
To evaluate cross-domain transfer Lai et al. (2022b); Liu et al. (2023a), we extract domain-labeled subsets from TED2025 according to the taxonomy of the SIB-200 benchmark. Instruction tuning is then performed on each domain using the MTC objective, with the resulting models evaluated across all other domains within SIB-200. Figure 8 illustrates the transfer performance.
Figure 8: Cross-domain generalization performance of instruction-tuned models using multi-way parallel data. Models trained on one domain are evaluated on all other domains from the SIB-200 benchmark.
Overall, instruction tuning with multi-way parallel data significantly improves domain transfer. The rich cross-lingual and cross-domain signals allow the model to learn domain-invariant features, enhancing robustness when confronted with novel topics and linguistic contexts. However, transfer performance remains limited in domains such as politics, sports, travel, and geography. We hypothesize that the high topical diversity, coupled with the relative sparsity of domain-specific examples in the training data, hinders the model’s ability to capture specialized patterns. Overcoming these limitations may require more balanced domain coverage or targeted data augmentation strategies.
7 Conclusion
In this paper, we construct a large-scale, high-quality multi-way parallel dataset covering 113 languages, with a maximum parallel degree of 50. This dataset provides a strong foundation for investigating the multilingual adaptation of LLMs. Using this dataset, we systematically explore best practices for adapting LLMs to multilingual tasks via multi-way parallel data. Our experiments reveal that multi-way data offers substantial advantages for both continued pretraining and instruction tuning, resulting in improved cross-lingual and cross-domain generalization. We further analyze key factors influencing model performance, including the degree of parallelism, the language combination strategies, and instruction training objectives.
Limitations
This work has the following limitations: (i) Due to limited computational resources, we employed parameter-efficient fine-tuning (PEFT) methods (LoRA) instead of full-parameter fine-tuning. While recent studies have demonstrated that LoRA achieves performance comparable to full fine-tuning across various tasks, our conclusions may still benefit from validation under full fine-tuning or alternative PEFT methods such as adapters or prefix tuning. (ii) Although our constructed dataset surpasses existing multi-way parallel corpora in both language coverage and maximum parallel degree, its overall size remains modest compared to large-scale unaligned multilingual datasets. To fully unlock the potential of multi-way parallel data for LLM adaptation, future work will focus on scaling up the dataset to further enhance multilingual performance.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grants No. 62236011 and No. T2341003), and by a grant from the Guoqiang Institute, Tsinghua University. Additionally, support was received from the European Research Council (ERC) under the European Union’s Horizon Europe Research and Innovation Programme (Grant Agreement No. 101113091), as well as from the German Research Foundation (DFG; Grant FR 2829/7-1).
Appendix A Statistics of the Constructed TED2025
In Section 3, we introduce TED2025 dataset and present the distribution of sentence counts across languages, along with the overall degree of parallelism (Figure 1) and translation quality compared with other multi-way corpora (Figure 2). To provide a more comprehensive overview, we further analyze domain coverage, fine-grained variations in parallelism, and the distribution of bitexts with respect to both quantity and quality.
Domain Coverage.
The TED2025 dataset is a multi-domain, multi-way parallel corpus encompassing 352 domains, with domain labels derived from TED Talks. Table 11 presents statistics on the number of talks per domain. We observe that the domains of global issues, education, technology, art, and business constitute the top five in terms of talk count. This statistical overview offers valuable insights into the dataset’s structure and facilitates a deeper understanding of its composition. Moreover, the domain labels enhance the dataset’s suitability for cross-lingual text classification tasks, positioning TED2025 as a robust benchmark for multilingual text classification.
Figure 9: Fine-grained parallelism in TED2025 dataset by tuple size.
Fine-grained Parallelism.
In Figure 1, we present an approximate count (ratio) of parallelism across different languages. For a more intuitive understanding of fine-grained parallelism in the TED2025 dataset, Figure 9 shows the tuple size corresponding to parallelism. Note that we count tuples rather than individual sentences. For instance, in a combination of six languages—English, French, Spanish, Russian, Arabic, and Chinese—a large tuple (en, fr, es, ru, ar, zh) encompasses all possible language combinations. The total number of such combinations is given by $C_{6}^{1}+C_{6}^{2}+C_{6}^{3}+C_{6}^{4}+C_{6}^{5}+C_{6}^{6}=6+15+20+15+6+1=63$ In contrast, the corresponding sentence count for this tuple is $6\times 1+15\times 2+20\times 3+15\times 4+6\times 5+1\times 6=192$ sentences. This tuple-based counting method offers a more precise analysis of parallelism in the dataset.
Quantity and Quality of Bitext.
In Figure 2, we compare the translation quality of TED2025 with existing multi-way parallel datasets, demonstrating the overall effectiveness of TED2025 translations. To offer a more comprehensive and intuitive view of translation quality, Table 1213141516, and 17 report COMET-QE score for all 4,765 language pairs included in the dataset. This analysis highlights that, despite being a multi-way parallel corpus, TED2025 can be readily decomposed into bilingual sentence pairs, making it suitable for training machine translation models or fine-tuning LLMs on specific language pairs. By providing both the number of bilingual pairs and their associated translation quality, we aim to support researchers in selecting suitable data for their specific translation tasks and in optimizing performance on targeted language pairs.
Appendix B Details of Experimental Setup.
LoRA Training
rank 8 batch size 8
alpha 32 learning rate 1e-04
dropout 0.1 lr schedule cosine
target all warmup ratio 0.1
Table 8: Hyper-parameters for continued pretraining and instruction tuning using LLaMA-Factory.
Training and Inference Setup.
Due to computational resource constraints, we adopt LoRA Hu et al. (2022) for continued pretraining and instruction fine-tuning of LLMs. All experiments are conducted using the LLaMA-Factory platform [^8] https://github.com/hiyouga/LLaMA-Factory Zheng et al. (2024), with training hyperparameters detailed in Table 8. Each experiment is run on 8 NVIDIA A100 (80GB) GPUs. For inference, we utilize the vLLM toolkit [^9] https://github.com/vllm-project/vllm, and the prompt templates used for each benchmark are partially sourced from PromptSource [^10] https://github.com/bigscience-workshop/promptsource as well as the respective original papers.
Evaluation Benchmarks.
We evaluate the trained model across a diverse set of tasks to comprehensively assess its capabilities in natural language understanding, commonsense reasoning, text generation, instruction following, and text classification. These tasks span multiple languages and domains, enabling us to measure both general and multilingual performance. Below, we outline the benchmarks used for each evaluation category:
  • Natural Language Understanding: We use the MMMLU benchmark, a multilingual extension of the widely adopted MMLU dataset Hendrycks et al. (2020), designed for evaluating the multitask language understanding abilities of large language models. MMMLU covers 14 languages and includes questions from a wide range of domains. We report accuracy across all tasks to measure performance.
  • Commonsense Reasoning: We evaluate the model using the XCOPA dataset Ponti et al. (2020). XCOPA tests a model’s ability to perform causal commonsense reasoning in multiple languages. The task involves selecting the most plausible cause or effect of a given premise from two alternatives, thus requiring both language understanding and reasoning skills.
  • Text Generation: We assess multilingual text generation performance using two benchmarks. First, FLORES-101 Goyal et al. (2022) is used to evaluate the model’s general generation quality across a broad set of high- and low-resource languages. Second, FLORES-200 Costa-Jussà et al. (2022) is employed to test zero-shot cross-lingual transfer capabilities—specifically, the model’s ability to generate high-quality outputs in languages it was not directly trained on.
  • Instruction Following: We use the multilingual variant of the IFEval benchmark Zhou et al. (2023), implemented by the Benchmax framework Huang et al. (2025), to evaluate the model’s ability to follow human instructions across diverse languages and task types. This benchmark focuses on the alignment between user instructions and model responses, which is critical for real-world applications of instruction-tuned models.
  • Text Classification: For evaluating domain-robust classification performance, we adopt the SIB-200 benchmark Adelani et al. (2024), which contains text classification tasks across 200 languages and multiple domains. This benchmark is particularly suited for testing the generalization and robustness of instruction-tuned models in a multilingual setting.
sk-ta $k$-ko eo-sw ja-zh nl-tr
BLEU COMET BLEU COMET BLEU COMET BLEU COMET BLEU COMET
Baseline 3.14 51.24 3.11 50.53 2.93 48.56 5.02 54.71 3.63 52.35
Unaligned 4.16 52.75 3.82 51.97 3.46 49.73 6.39 55.84 4.75 52.27
Multi-Way 6.31 53.25 5.38 52.66 5.14 51.35 8.06 56.27 6.82 53.50
Table 9: Translation performance of models trained on aligned, unaligned, and multi-way data setups for five non-English-centric language pairs.
Appendix C Non-English-Centric Translation Performance
In addition to the primary experiments on English–X translation directions presented in Section 4.1, we also investigate the performance of our models on non-English-centric translation pairs. This analysis is motivated by the potential of multilingual models to perform well in settings where neither language involved is English. To this end, we randomly select five language pairs (sk-ta, kk-ko, eo-sw, ja-zh, and nl-tr) that do not involve English as either the source or target language. We then evaluate the translation performance of the pretrained models under three different data configurations: Baseline, Unaligned, and Multi-way, as described in Section 4.1. The results, including BLEU and COMET scores, are presented in Table 9.
We observe that: (1) Models trained on aligned data consistently outperform those trained on unaligned data across all translation pairs, indicating that aligned corpora are essential for achieving high translation quality, regardless of whether English is involved. (2) Models trained on unaligned data show a significant improvement over the baseline, suggesting that even in the absence of high-quality aligned corpora, leveraging unaligned multilingual data can still enhance translation performance. (3) The multi-way approach consistently yields superior results compared to both the baseline and unaligned models. This highlights the advantages of multi-way training for multilingual translation tasks, as it not only improves performance on English-centric tasks but also significantly boosts translation quality for non-English-centric language pairs.
Figure 10: Representation alignment with instruction tuning across different training objectives.
English Pivoting Combinations
Group Language List
group 1 en,vi,ar,bg,de,es,fr,he,it,ja,ko
group 2 en,nl,pl,pt-br,zh-cn,ar,bg,de,es,fr,he
group 3 en,el,vi,ar,bg,de,es,fr,he,it,ja
group 4 en,el,ar,bg,de,es,fr,he,it,ja,ko
group 5 en,fa,hu,ar,bg,de,es,fr,he,it,ja
Language Family Combinations
Group Language List Language Family
group 1 bg,pl,ru,sr,uk Slavic
group 2 el,it,kmr,sq,tr Indo-European
group 3 en,ko,my,sq,zh-tw Indo-European (en, sq), Sino-Tibetan (zh-tw), Koreanic (ko)
group 4 fr,hu,hy,lv,vi Indo-European (fr, hy), Uralic (hu, lv), Austroasiatic (vi)
Table 10: Language configurations for different sampling groups.
Appendix D Configuration of Different Language Combinations
As discussed in Section 5.3, we define four distinct language groups for our analysis. The specific configurations of these groups are presented in Table 10.
Appendix E Representation Alignment with Instruction Tuning
In this section, we explore how each tuning objective reshapes the model’s internal multilingual embeddings, using the same alignment metrics as in Section 4.3. Figure 10 shows that MT-tuned models achieve the highest SVCCA score, indicating tighter alignment among semantically equivalent sentences across languages. In contrast, CLTS yields minimal alignment gains despite its similarity focus, likely because binary similarity labels lack the contextual richness of translation pairs.
Topic #Talks Topic #Talks Topic #Talks Topic #Talks Topic #Talks
global issues 10334 race 154 coronavirus 70 inclusion 36 resources 18
education 9126 writing 154 sex 69 solar system 35 paleontology 17
technology 6323 physics 153 pandemic 69 sight 35 dinosaurs 17
art $58$ gender 144 product design 68 Audacious Project 35 television 17
business 5694 exploration 143 algorithm 68 industrial design 34 exercise 17
Life 5623 neuroscience 142 aging 67 sound 34 blockchain 17
health 5561 family 140 empathy 67 behavioral economics 34 fungi 16
science 4819 architecture 138 justice system 67 online privacy 34 public speaking 16
entertainment 4717 emotions 137 goals 66 travel 33 nuclear energy 16
design 2313 humor 136 urban planning 65 magic 33 PTSD 16
Humanities 1048 happiness 132 crime 65 trees 33 driverless cars 16
TED-Ed 941 universe 131 machine learning 65 weather 33 bees 15
animation 879 AI 131 compassion 64 public space 33 geology 15
culture 847 illness 131 investing 62 TED Prize 32 smell 15
social change 783 entrepreneur 130 fish 62 bioethics 31 Alzheimer’s 15
TEDx 734 decision-making 130 demo 62 military 30 shopping 15
society 712 language 130 disability 62 illusion 30 Autism spectrum disorder 15
history 617 self 130 conservation 62 human rights 29 vulnerability 15
innovation 518 photography 127 feminism 61 theater 28 TED Connects 15
humanity 500 policy 126 Planets 60 solar energy 28 toys 14
biology 482 religion 124 renewable energy 57 biosphere 28 Antarctica 14
communication 453 media 121 code 57 maps 27 science fiction 14
future 434 democracy 120 women in business 57 spoken word 27 Islam 14
creativity 413 india 120 chemistry 57 heart 27 neurology 14
climate change 412 energy 118 Middle East 57 Slavery 27 Moon 13
community 407 motivation 115 immigration 57 sexual violence 27 3D printing 13
personal growth 395 philosophy 114 Best of the Web 56 surveillance 27 body language 13
environment 392 film 113 natural disaster 55 manufacturing 27 hearing 13
activism 371 violence 113 dance 55 trust 27 meditation 13
sustainability 341 literature 112 consciousness 54 archaeology 26 Brazil 12
performance 325 parenting 111 marine biology 54 AIDS 25 graphic design 12
psychology 325 journalism 111 virus 53 virtual reality 25 ebola 12
medicine 323 money 111 statistics 52 gardening 25 suicide 12
brain 323 potential 110 depression 52 nanotechnology 25 wind energy 12
music 316 ancient world 109 microbiology 52 Europe 25 coral reefs 12
work 316 social media 108 china 51 biomimicry 24 rivers 12
economics 308 love 107 electricity 51 drones 24 international relations 12
health care 307 poetry 107 plants 51 mindfulness 24 glaciers 12
nature 300 law 107 Vaccines 51 quantum 24 augmented reality 12
collaboration 288 pollution 106 Asia 49 aliens 24 worklife 12
politics 279 biodiversity 105 ethics 49 friendship 24 rocket science 11
animals 271 software 103 bacteria 48 encryption 24 Christianity 11
women 269 visualizations 103 farming 48 medical imaging 24 Sun 11
TED Fellows 267 teaching 100 prison 47 South America 23 bullying 11
human body 267 international development 99 refugees 47 telescopes 23 CRISPR 11
storytelling 259 finance 99 TED en Español 47 birds 23 String theory 10
invention 250 genetics 96 gaming 46 Big Bang 22 homelessness 10
identity 245 death 96 fear 46 Mission Blue 22 grammar 9
kids 243 books 94 natural resources 46 Surgery 22 typography 8
engineering 241 TED Residency 94 LGBTQIA+ 46 protest 22 Buddhism 8
leadership 234 biotech 92 terrorism 44 painting 22 asteroid 8
government 233 beauty 92 philanthropy 44 interview 21 deextinction 8
equality 223 work-life balance 89 personality 44 library 21 cryptocurrency 8
public health 219 robots 88 marketing 43 plastic 21 metaverse 8
medical research 218 poverty 87 drugs 43 Mars 21 conducting 7
Internet 214 water 84 sociology 43 Egypt 21 bionics 7
computers 201 transportation 83 curiosity 43 pregnancy 21 NASA 7
Africa 197 cancer 83 TEDMED 43 synthetic biology 21 atheism 5
data 194 astronomy 82 indigenous peoples 42 Transgender 21 pain 5
cities 193 agriculture 82 sleep 41 prosthetics 20 forensics 4
disease 192 DNA 79 diversity 40 museums 20 microbes 4
space 187 success 79 fashion 40 addiction 20 NFTs 4
war 183 ecology 77 discovery 40 TED Membership 19 Judaism 3
United States 181 infrastructure 77 corruption 39 primates 18 street art 3
food 174 cognitive science 76 time 39 cyber security 18 Hinduism 1
math 173 sports 74 consumerism 38 astrobiology 18 reproductive health 1
ocean 168 youth 73 productivity 38 dark matter 18 crowdsourcing 1
Countdown 168 comedy 73 Anthropocene 38 botany 18 veganism 1
evolution 162 insects 71 capitalism 38 fossil fuels 18
mental health 159 memory 70 flight 36 blindness 18
relationships 154 anthropology 70 UX design 36 TED Books 18
Table 11: Statistics on the number of talks per domain in the TED2025 dataset.
Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE
ar_en 1242114 74.58 ko_nl 603056 61.26 en_id 414970 78.6 bg_cs 290870 72.28 fa_sq 144330 62.35 hi_ja 101055 64.02 ja_lv 77169 59.48
en_es 1171895 78.54 nl_ru 602559 67.92 ar_uk 414459 63.47 de_uk 290812 70.01 pt-br_sq 143939 65.7 ca_fa 100797 65.67 ckb_ru 77119 66.2
ar_es 1076362 67.49 fa_fr 601008 66.3 el_nl 411252 71.5 cs_ro 289965 66.43 pl_sq 143091 58.49 lt_th 100559 64.29 hy_it 77029 72.39
ar_zh-cn 994593 56.73 hu_ru 599111 66.95 el_ro 410986 72.64 ro_th 289049 68.12 sk_th 142947 67.93 kmr_vi 99796 57.87 id_kmr 76830 68.53
en_zh-cn 973667 68.84 fa_zh-cn 597792 58.76 hr_zh-tw 409256 65.99 pt_tr 288876 65 da_he 142920 68.94 my_ro 99349 59.51 lv_pt-br 76810 64.23
en_fr 942555 78.07 en_pl 597234 71.96 hr_ko 406822 64.74 id_nl 288406 70.38 es_lt 142342 66.23 da_uk 99161 71.01 hu_lv 76732 61.74
ar_fr 930909 65.64 es_hu 595397 67.29 el_fa 405959 65.36 it_pt 286615 70.47 da_ja 142244 60.65 my_pl 98672 64.29 es_hy 76522 72.01
ar_ko 922379 59.11 fa_tr 595201 63.42 hr_ru 403875 70.08 de_id 286157 68.35 pt_th 142194 68.07 ca_en 97477 76.76 he_hy 76382 65.64
es_zh-cn 917132 61.02 it_ro 594433 69.53 bg_hu 402557 67.52 el_uk 284286 69.11 he_lt 141827 65.44 ca_pl 96318 65.14 fr_hy 76335 71.77
en_pt-br 916532 76.85 ja_pl 592311 60.64 id_zh-cn 402550 63.86 pt_vi 283273 68.58 sv_uk 141678 71.52 hi_hu 95946 69.46 bg_mk 76211 66.44
ar_zh-tw 904984 58.14 pt-br_ro 591880 69.36 uk_zh-tw 402129 59.77 bg_uk 280180 66.54 lt_zh-cn 140519 58.59 he_hi 95463 65.91 fa_sl 75967 67.71
es_fr 891473 70.34 pl_pt-br 587885 64.03 es_id 400666 69.39 sr_th 274803 67.47 da_pt-br 139360 70.55 id_lt 95440 64.5 ckb_fa 75797 58.61
en_ko 890124 70.57 he_ro 587631 67.31 en_pt 400349 77.17 el_id 271766 70.96 hu_sq 139156 59.85 da_th 95366 68.44 hy_zh-cn 75783 61.69
ar_pt-br 879212 67.73 pl_tr 586000 61.75 el_hu 399729 68.3 de_th 270867 65.05 da_tr 139059 63.87 kmr_pl 94413 65.86 sl_vi 75423 66.19
en_it 878170 77.63 fr_hu 585533 67.48 ru_uk 398976 66.93 cs_el 269611 71.02 lt_tr 138684 60.07 da_id 94277 69.27 hu_sl 75210 64.94
ar_ru 877084 62.75 es_nl 583927 68.53 ko_uk 397906 60.44 fa_pt 266363 69.15 sv_th 138653 70.03 ar_fi 93954 63.87 pt-br_sl 75095 67.73
en_zh-tw 875551 70.77 ro_tr 583912 62.87 hu_sr 395571 68.03 ja_pt 264092 61.77 fr_lt 138611 66.97 id_my 93952 63.59 hr_my 74708 65.31
es_pt-br 852199 68.21 fr_nl 581625 68.29 en_uk 395281 74.99 el_th 263718 65.59 lt_vi 137996 64.4 kmr_ro 93814 63.76 hy_pl 74621 68.75
en_ru 850952 73.92 it_nl 579681 69.3 es_hr 394545 72.79 bg_id 262916 71.63 es_my 137828 62.85 ca_nl 93527 68.23 sk_sq 74607 59.87
ar_it 847883 68.63 ar_de 579476 63.12 he_hr 393849 71.16 hr_uk 259520 69.92 nl_sq 137756 65.76 hi_vi 93155 65.35 ca_id 74599 66.27
es_ko 840427 63.57 fa_ja 577117 60.12 pl_sr 392935 65.49 bg_th 259002 63 da_vi 137054 69.4 ar_mk 92840 63.44 de_my 74489 61.94
ko_zh-tw 836982 57.84 ja_ro 577022 58.02 de_el 392903 69.08 cs_sr 251810 67.23 ja_lt 136887 59.28 ckb_zh-cn 92736 55.79 hy_pt-br 74342 71.54
ar_vi 835188 60.61 en_nl 574465 74.32 hr_tr 391818 64.98 he_pt 251208 69.61 ko_my 136833 62.55 ca_de 92300 65.77 hy_vi 74020 68.65
ar_tr 834805 57.72 nl_zh-cn 573272 62.01 bg_fa 391024 63.69 cs_hr 246496 68.76 de_sq 136631 62.52 fi_zh-tw 92202 62.77 el_hi 74002 67.9
es_zh-tw 826342 62.19 fa_it 569719 69.54 hr_it 389151 71.47 id_sr 244934 70.81 cs_sv 136553 69.84 mk_zh-tw 91924 60.64 pl_sl 73913 63.02
ko_zh-cn 820514 55.92 he_nl 569437 67.38 ar_th 387921 60.43 th_uk 240998 65.41 my_zh-cn 136213 54.99 ko_mk 91599 58.57 mk_ro 73468 71.39
en_vi 819601 74.7 de_ko 566540 60.88 fr_hr 387881 72.49 hr_th 233343 69.03 bg_sq 136032 64.2 fi_ko 91568 62.31 bg_lv 73230 66.62
ko_ru 817529 58.43 de_ru 566165 66.03 en_hr 387342 80.44 hu_pt 231917 66.73 en_sq 135269 75.78 fi_ru 91544 67.73 fa_lv 73228 65.09
zh-cn_zh-tw 814894 59.22 de_zh-tw 565984 60.88 id_ko 386648 63.97 hr_id 225917 69.17 my_ru 134946 57.26 my_sr 90728 63.79 kmr_uk 73167 63.34
fr_ko 812164 61.91 hu_tr 565854 61.79 hr_zh-cn 386282 63.07 ar_sv 225106 68.86 hr_pt 134940 69.73 ckb_es 90328 66.87 hu_hy 73131 68.61
ru_zh-tw 808709 57.9 fa_pt-br 563585 68.58 nl_sr 383972 70.94 cs_uk 224284 70.41 fa_lt 133962 64.81 sk_sv 90046 67.92 fi_ro 73106 65.46
fr_zh-tw 807365 61.74 nl_pt-br 562777 69.25 ro_sr 383556 69.62 ar_sk 219162 66.47 lt_pt-br 133915 64.4 kmr_sr 89752 63.3 de_lv 72910 67.18
en_tr 803643 69.44 hu_zh-cn 562549 61.05 fr_id 380333 69.33 ru_sk 216633 67.31 da_pl 133738 67.33 bg_ca 89729 67.71 nl_sl 72893 69.34
fr_zh-cn 802960 59.75 fa_he 561511 63 id_zh-tw 380316 65.19 pt_ro 216447 70.24 da_fa 133336 67.54 fi_tr 89546 63.01 my_pt 72689 62.59
es_it 802859 71.06 pl_vi 559884 63.99 ja_uk 380238 60.28 ko_sk 216434 60.12 id_sk 132944 68.69 mk_ru 89477 64.36 de_hi 72551 68.11
es_ru 801560 67.16 hu_pt-br 559813 66.58 hr_ja 378739 62.86 sk_zh-tw 215458 63.18 da_nl 132314 68.62 fi_zh-cn 89334 62.16 lv_nl 72451 66.67
fr_ru 792541 66.25 ja_nl 558716 60.73 ar_pt 378332 68.92 ru_sv 214353 67.74 lt_pl 132181 64.39 ca_ro 88907 66.49 hy_ja 72415 62.09
ar_ja 784497 56.38 nl_tr 557514 64.76 it_uk 377321 70.33 cs_th 213428 67.02 my_tr 131897 63.15 es_fi 88720 67.55 de_hy 72367 70.66
fr_pt-br 782057 69.2 de_fr 555403 67.25 tr_uk 376704 62.37 sv_zh-tw 211023 65.56 bg_da 131380 71.93 mk_tr 88567 60.92 bg_hy 72117 70.43
ru_zh-cn 782001 56.46 de_es 555297 67.4 es_uk 376019 70.15 it_sk 210780 69.94 da_hu 130980 68.1 it_mk 88419 72.46 ar_gl 71993 64.86
es_vi 773676 66.34 de_zh-cn 550745 59.68 hr_pt-br 375441 70 ko_sv 210646 61.02 my_zh-tw 130701 58.12 es_mk 88223 70.91 hi_sr 71827 72.17
pt-br_zh-cn 771080 61.87 de_it 548882 68.86 fr_uk 374373 67.84 he_sk 209261 68.5 hu_lt 130019 63.01 fr_mk 87857 69.59 lt_sk 71823 63.62
ko_pt-br 769640 60.88 hu_ja 548815 60.43 id_ru 373973 68.57 id_uk 208434 70.49 el_sq 130019 67.52 mk_zh-cn 87776 58.56 lt_sv 71797 66.13
ar_he 769088 63.17 hu_it 546990 68.89 he_uk 373946 66.87 es_sk 207954 70.01 da_de 129984 67.71 fi_it 87350 68.87 hy_nl 71760 71.78
it_zh-cn 768684 62.08 fa_vi 540771 62.78 ru_th 371985 62.57 fr_sk 207151 70.04 fa_my 129218 61.56 kmr_nl 87266 65.77 el_fi 71717 67.66
en_ja 767383 67.19 de_pt-br 539297 68 bg_el 370591 71.05 id_th 206594 70.04 lt_nl 129116 65.92 ca_el 87260 71.04 en_gl 71708 76.4
pt-br_zh-tw 765142 63.79 ro_vi 538958 66.99 id_tr 370340 64.68 en_sv 206079 78.4 ar_kmr 129089 58.57 fi_he 87220 64.39 da_sk 71598 69.62
it_ko 763728 63.57 pl_ro 538936 64.87 th_zh-tw 370162 62.54 es_sv 204734 70.93 ro_sq 129001 62.8 ckb_tr 87019 55.63 el_mk 71404 70.38
ko_tr 762690 62.69 de_he 538274 65.08 ko_th 369872 59.23 sk_zh-cn 204644 61.05 en_lt 128870 77.27 fi_fr 86735 68.32 kmr_th 71125 55
fr_it 759270 70.8 he_hu 537399 65.4 hr_vi 367480 69.78 fr_sv 204374 70.21 id_sv 128754 71.48 ca_sr 86525 69.29 fa_hy 70864 68.08
it_zh-tw 758320 64.39 ar_bg 535507 64.18 uk_zh-cn 365583 58.46 pl_pt 203742 65.16 ar_hi 128727 61.94 lv_ru 85431 66.65 hi_id 70512 72.84
vi_zh-cn 758118 58.82 bg_zh-tw 529405 60.27 id_vi 365422 69.2 cs_id 203628 67.15 hr_sq 127983 61.56 ar_lv 85384 64.03 ar_fr-ca 70482 65.28
es_tr 758036 63.25 bg_ko 529377 60.53 es_pt 364577 69.57 ja_sk 202798 61.4 da_en 127603 77.74 fi_ja 85205 61.57 fi_hr 70341 67.46
it_ru 755860 69.14 bg_ru 527966 64.19 uk_vi 363859 63.49 sv_tr 202544 65.03 en_hi 127575 74.11 lv_zh-tw 85107 61.77 en_sl 70278 76.52
tr_zh-tw 754441 60 nl_pl 526107 67.7 id_pt-br 361604 69.89 sk_tr 202354 63.07 fr_my 127214 61.5 he_mk 84836 69.53 bg_sl 70217 68.85
pt-br_ru 746806 67.94 nl_vi 524453 67.49 en_th 360968 74.44 it_sv 202096 71 my_vi 126920 58.12 el_kmr 84364 66.19 sq_sv 70151 65.62
ru_tr 743416 59.64 de_tr 522926 62.2 fa_uk 360247 65.44 sv_zh-cn 202046 63.58 en_kmr 126910 74.74 ar_sl 84275 66.2 pt_sk 69947 68.49
ja_ko 741597 63.41 de_en 521558 72.11 fa_hr 359517 71.19 sk_vi 200977 66.32 da_ro 124242 69.18 ko_lv 84272 58.64 de_sl 69903 68.57
ja_ru 738225 57.25 de_ja 520960 58.58 id_it 358047 70.92 he_sv 200731 69.67 de_lt 123263 67.47 hu_mk 84242 67.94 fr-ca_zh-tw 69524 60.2
ja_zh-tw 738006 56.72 bg_it 518477 72.39 he_th 356554 63.74 nl_pt 200282 70.49 sq_sr 122703 65.64 de_kmr 84191 66.03 fr-ca_it 69422 70.03
tr_zh-cn 735014 57.29 ar_el 516272 67.15 hr_pl 355076 66.9 pt-br_sk 199502 69.46 bg_lt 122347 69.67 ja_mk 83889 57.56 ckb_he 69176 64.45
he_ko 731416 58.54 bg_es 515032 71.15 ar_cs 353364 68.36 pl_sk 197076 64.89 he_my 122318 59.17 fa_fi 83877 65.84 es_fr-ca 69123 70.11
fr_tr 730265 62.45 bg_zh-cn 513676 57.78 pt-br_uk 352163 71.05 nl_sk 195616 69.98 kmr_zh-tw 122148 57.91 en_lv 83504 76.43 el_sl 69016 69.84
en_he 727009 77.91 bg_he 513379 67.16 it_th 352152 68.46 fa_sv 195104 68.61 cs_pt 121276 65.99 fi_pt-br 83408 66.91 mk_sr 69016 72.21
he_ru 725834 63.89 bg_fr 512323 70.2 he_id 351414 68.37 ja_sv 195050 61.82 kmr_ko 120297 61.03 fi_hu 83367 64.89 fr-ca_ru 68995 66.06
he_zh-tw 724228 58.85 de_pl 511972 66.79 ja_th 351200 59.68 pt-br_sv 193655 70.9 ja_my 120050 62.06 my_th 83330 58.22 lv_ro 68727 60.9
fr_ja 721692 58.11 fa_hu 507281 66.84 hr_hu 349614 67.19 sv_vi 192166 69.76 kmr_ru 118804 64.93 sl_zh-tw 83227 62.43 fr_fr-ca 68723 69.2
es_ja 718188 59.91 nl_ro 506646 69.14 th_zh-cn 348388 60.81 fa_sk 191244 69.65 kmr_tr 118198 54.63 fi_vi 83022 66.27 hi_hr 68674 74.16
it_pt-br 715066 68.82 hu_vi 503814 64.36 cs_ru 347890 67.99 hu_sk 190231 66.59 my_pt-br 117759 62.18 el_my 82977 60.53 ckb_ja 68626 59.56
ko_vi 710975 59.13 ar_sr 502347 68.59 th_vi 346956 65.55 en_sk 189541 77.09 fr_kmr 117606 65.11 ru_sl 82964 66.75 fr-ca_ko 68476 59.36
es_he 707993 68.38 el_ru 500339 67.92 es_th 346791 66.67 hu_sv 188563 67.95 it_my 117321 65.07 ca_hr 82861 69.6 hr_mk 68455 70.18
ja_zh-cn 705576 54.78 el_ko 498849 61.49 hr_nl 346217 70.46 pt_sr 186440 71.36 hi_ko 117013 64.5 hi_ro 82802 71.13 fr-ca_he 68153 65.95
pt-br_tr 700606 64.28 bg_pt-br 495814 72.04 cs_zh-tw 346046 62.56 el_pt 186362 72.91 ckb_en 116963 74.39 es_lv 82169 64.47 fr-ca_pl 68105 67.64
he_zh-cn 697132 57.13 de_vi 495041 64.27 de_sr 345374 69.83 de_sk 185010 69.5 cs_sq 116017 60.89 ko_sl 82129 60.3 fr-ca_zh-cn 68015 58.74
fr_he 695231 66.42 el_zh-tw 494332 62.01 cs_ko 344750 60.34 bg_sk 184233 73.08 lt_ro 115112 62.5 fi_pl 81916 65.53 fi_sr 67854 67.87
it_tr 690366 64.47 el_en 493620 77.66 th_tr 344740 60.06 nl_sv 180944 70.32 lt_sr 115048 64.39 mk_pt-br 81833 72.54 ar_nb 67649 68.08
ru_vi 690178 60.95 bg_pl 492192 69.55 fr_th 343311 65.25 pl_sv 178707 66.06 hi_zh-tw 114738 65.38 it_lv 81771 65.67 fr-ca_pt-br 67093 68.83
vi_zh-tw 688560 61.05 el_es 490555 71.36 id_ja 343105 63.25 ro_sk 176638 68.43 es_hi 114569 68.86 fa_mk 81666 65.49 ckb_id 67079 66.97
it_ja 686850 60.84 bg_ja 489386 56.99 cs_it 340789 67.32 de_pt 174625 68.77 ar_ca 114566 63.29 fi_nl 81595 68.47 hr_lv 67057 63.75
he_it 686200 69.73 en_sr 487888 79.1 cs_es 336661 67.56 el_sk 171053 72.61 da_el 114431 72.86 ckb_fr 81115 63.4 fr-ca_vi 66872 65.72
fr_vi 685777 65.31 de_nl 486354 66.88 hr_ro 334834 70.26 en_my 170208 73.07 el_lt 114330 68.74 pt_sv 81032 71.2 da_sv 66803 71.18
it_vi 685674 68.08 de_ro 486226 69.47 el_sr 334743 72.97 id_pt 169049 69.77 hr_lt 114172 65.4 my_uk 80952 59.6 bg_fr-ca 66506 70.99
ar_fa 682164 60.73 bg_tr 485945 60.13 hu_uk 333811 68.12 ro_sv 164123 68.74 sq_uk 113176 64.65 bg_fi 80834 68.46 el_lv 66307 67.2
ja_tr 680787 61.08 el_fr 485575 71.74 cs_he 332986 65.89 hr_sk 162606 69.15 fa_kmr 113100 59.6 mk_vi 80774 65.39 fr-ca_tr 66157 61
he_ja 673308 55.62 sr_zh-tw 482580 65.34 cs_fr 332977 68.71 sk_sr 160805 67.64 es_kmr 113087 68.1 mk_pl 80386 69.76 es_gl 66101 66.31
ja_pt-br 671352 59.75 fa_ro 479887 68.18 fa_id 332662 68.99 de_sv 160635 69.15 fr_hi 112156 67.53 fr_lv 80379 65.26 hr_hy 65865 72.01
ar_ro 668130 68.34 el_zh-cn 479001 60.39 cs_tr 331127 62.75 ar_my 160508 57.53 ca_zh-tw 112027 61.01 es_sl 80372 68.14 fr-ca_ja 65632 57.49
en_fa 664436 75.19 ko_sr 478944 63 fa_th 330735 62.57 el_sv 159937 72.99 he_kmr 111431 65.07 he_lv 80147 64.29 gl_ko 65577 62.9
he_tr 663973 62.41 ru_sr 478933 67.36 pl_uk 330051 67.26 ar_sq 159140 62.05 ja_kmr 111371 59.52 hi_pl 80095 71.04 cs_fi 65511 64.03
he_pt-br 662966 68.85 fa_pl 476653 67.33 cs_zh-cn 329173 61.13 sr_sv 158836 70.71 ca_ko 110398 60.57 lv_zh-cn 79985 60.06 hr_sl 65114 68.37
ar_pl 654671 64.63 hu_ro 474206 66.37 pt_zh-cn 329068 64.55 bg_sv 156445 74.01 ca_tr 110139 60.04 he_sl 79961 67.35 gl_zh-tw 65095 63.25
tr_vi 648264 57.77 bg_en 473821 78.53 fr_pt 326890 69.68 ko_sq 155951 55.95 da_sr 110097 71.29 my_nl 79943 63.53 es_nb 65092 70.81
ar_hu 647420 66.11 el_it 473609 73.42 bg_hr 326692 73.02 ru_sq 155867 60.9 da_hr 109867 72.08 ckb_zh-tw 79830 58.89 de_fr-ca 65026 68.09
pt-br_vi 646236 67.56 hu_nl 470590 67.11 cs_ja 326516 58.71 sq_zh-tw 155482 59.27 hi_ru 109417 66.55 lv_tr 79708 61.08 fr-ca_nl 64973 69.54
he_vi 636007 64.84 hu_pl 469374 65.04 nl_uk 326003 70.6 pt_uk 153176 70.84 kmr_zh-cn 109247 54.3 bg_kmr 79684 61.87 el_hy 64914 68.75
ro_zh-tw 635819 62.38 bg_nl 468125 70.5 de_hr 325841 71.35 sq_zh-cn 152895 56.39 ca_es 109050 69.31 hr_kmr 79446 66.96 nb_ru 64852 69.77
fa_ko 635023 61.51 fa_nl 465950 68.18 bg_sr 325120 70.8 ar_da 152630 67 ar_ckb 108785 60.49 en_mk 79374 77.72 cs_mk 64817 67.9
fa_zh-tw 634299 60.38 el_he 463842 67.15 cs_vi 324934 67.25 es_sq 152545 64.74 ca_ru 108220 65.09 lv_vi 79339 65.21 bg_hi 64597 66.02
ko_ro 633363 59.49 el_ja 463791 59.36 pt-br_th 324124 67.89 cs_sk 152528 68.33 lt_uk 108111 66.67 ca_cs 79322 68.04 lv_sr 64173 65.19
ko_pl 633060 59.97 bg_ro 463670 71.81 ko_pt 319151 62.46 it_sq 152509 66.18 hi_tr 107996 64.89 sl_tr 79270 61.82 ko_nb 64095 61.56
pl_ru 632873 65.29 el_pt-br 463625 72.6 cs_pt-br 318949 65.77 ar_lt 152127 64.79 ca_fr 107286 67.65 ckb_pt-br 79170 68.37 bg_my 63905 58.25
pl_zh-tw 631594 60.58 bg_vi 462776 63.62 cs_en 317213 74.91 da_ko 150697 62.11 id_sq 107217 64.18 it_sl 79156 68.16 en_hy 63894 79.32
en_ro 631593 79.14 es_sr 461004 71.37 id_pl 316519 66.68 da_zh-tw 150571 63.76 hu_my 106915 62.43 de_mk 78930 68.88 cs_kmr 63711 65.99
en_hu 629406 75.6 el_tr 460556 62.99 ro_uk 314492 69.88 fr_sq 150387 62.89 kmr_pt-br 106726 66.87 fr_sl 78904 67.74 mk_uk 63570 64.54
ja_vi 629331 57.89 bg_de 458941 67.67 pl_th 314468 66.08 da_ru 149665 68.57 hi_zh-cn 106417 63.57 en_fi 78738 77.43 hy_ro 63389 68.62
ro_ru 626305 68.47 ja_sr 454498 62.87 cs_pl 314358 65.68 sq_tr 148924 57.5 ca_it 106159 68.91 hy_ru 78701 68.99 fr_gl 63387 66.4
es_ro 625186 69.58 fr_sr 453800 71.17 hu_id 310989 65.83 he_sq 148816 63.26 fa_hi 106046 64.82 mk_nl 78690 72.62 ckb_kmr 63270 65.26
fa_ru 623287 63.37 sr_tr 449990 63.67 pt_zh-tw 310985 65.54 lt_zh-tw 148626 60.17 cs_lt 105656 63.84 ca_uk 78668 67.92 hi_uk 63175 68.17
ar_nl 621126 66.01 sr_zh-cn 449486 61.44 cs_nl 308643 66.66 sq_vi 148180 63.74 cs_da 105170 66.01 ckb_ko 78665 63.2 ckb_it 63157 68.93
es_pl 618290 67.63 it_sr 448764 71.37 id_ro 305376 69.65 hr_sv 147522 71.29 sq_th 104846 63.57 sl_zh-cn 78644 60.02 sl_sr 63125 68.73
ro_zh-cn 614544 60.73 he_sr 447876 70.13 hu_th 304845 65.07 sk_uk 147414 69.22 hu_kmr 104539 65.3 hi_nl 78635 72.04 gl_ru 63102 67.12
fr_ro 614455 69.17 el_vi 444442 67.1 cs_fa 303677 66.13 da_es 147382 70.09 ca_he 104482 65.44 ja_sl 78545 61.45 nb_zh-cn 62836 63.05
it_pl 613374 67.21 ar_id 430301 68.25 sr_uk 301848 69.18 bg_pt 146973 72.17 it_kmr 104349 69.05 hy_zh-tw 78447 63.17 it_nb 62824 70.35
es_fa 613256 67.5 de_hu 430188 67.48 cs_hu 297678 63.77 lt_ru 146607 66.31 ca_zh-cn 104326 59.69 de_fi 78395 66.89 cs_hy 62783 70.48
pl_zh-cn 611358 58.42 pt-br_sr 429125 70.82 pt_pt-br 297507 68.43 ko_lt 146350 58.05 ca_ja 102807 58.69 hy_ko 78284 64.25 sl_uk 62693 67.8
he_pl 611074 65.25 de_fa 424941 66.11 pt_ru 297277 69.66 da_fr 146113 70.52 ca_hu 102305 64.51 lv_pl 77852 63.41 nb_vi 62665 70.29
fr_pl 609128 67.84 el_pl 423889 69.18 hr_sr 295264 71.79 da_zh-cn 145736 61.59 hi_pt-br 102293 72.4 ar_hy 77713 66.9 nb_zh-tw 62665 64.33
hu_zh-tw 607791 61.76 sr_vi 423151 67.63 el_hr 295227 74.19 da_it 145604 71.11 ca_pt-br 102038 67.07 ca_th 77473 64.57 gl_tr 62607 63.06
nl_zh-tw 606375 64.74 ar_hr 421331 69.64 nl_th 294763 68.05 ja_sq 145482 54.99 ca_vi 101669 64.13 hy_tr 77387 65.43 fr_nb 62423 69.84
hu_ko 606236 62.08 fa_sr 419593 68.36 cs_de 294216 67.96 it_lt 144856 67.57 hi_it 101118 73.9 ckb_vi 77253 59.74 fa_fr-ca 62358 64.9
Table 12: COMET-QE score and the count of sentences for all 4,765 language pairs in TED2025 (part I).
Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE
ro_sl 62309 65.04 nb_th 41483 68.87 id_mr 28855 62.92 bn_vi 23238 66.1 en_sw 20023 77.74 mn_sk 17389 57.28 sk_ta 13162 62.77
fa_gl 62195 66.72 hy_sq 41172 66.65 de_mn 28786 55.66 az_ru 23136 58.27 it_kk 19975 64.07 hi_lv 17362 70.18 gu_uk 13126 67.71
hi_th 61766 63.11 mr_zh-tw 41051 59.5 hr_mr 28741 61.72 bs_tr 23103 64.35 kk_vi 19960 57.71 fi_my 17350 62.66 fr-ca_hi 13071 67.4
fr-ca_hu 61740 67.53 ca_pt 40944 67.44 ar_ta 28701 59.23 fr_zh 23059 55.3 kk_ru 19953 62.81 lt_mn 17323 50.67 fa_uz 12972 58.59
nb_tr 61721 64.88 ko_mr 40861 57.46 ru_ur 28670 66.47 es_zh 23042 55.89 eo_vi 19940 62.58 bn_el 17306 66.77 da_mn 12894 57.22
gl_zh-cn 61551 61.4 mr_ru 40852 61.09 bg_mr 28482 59.18 az_it 23039 62.1 be_fa 19909 61.3 hi_mk 17286 66.4 bs_sv 12879 71.79
ckb_ro 61414 65.67 ca_da 40750 64.95 mn_ro 28440 54.6 bs_fa 23027 68.83 lv_mk 19841 65.22 kk_pl 17265 64.42 he_uz 12862 56.45
hy_id 61346 69.19 sl_sv 40580 68.09 ca_hy 28369 69.32 bs_pt-br 22950 70.16 mr_sv 19837 65.29 be_zh-cn 17197 55.22 ko_uz 12825 57.94
lv_uk 61183 66.36 hr_ka 40504 67.28 pt_sl 28315 67.3 az_zh-tw 22930 57.4 bg_bn 19824 65.58 be_fr 17161 64.58 es_uz 12820 55.61
he_nb 61043 69.77 ar_et 40194 60.93 hi_kmr 28312 61.12 bs_it 22902 71.89 ms_th 19782 66.91 ka_lv 17126 63.77 eo_sv 12787 67.71
fi_th 60689 65.39 el_ka 40126 63.99 ar_ur 28260 63.75 nl_ur 22882 71.69 gu_it 19781 71.55 bn_uk 17121 66.27 be_de 12733 63.47
cs_lv 60607 61.5 nb_uk 39907 71.13 ru_ta 28237 59.77 nb_pt 22869 71.02 mn_sv 19765 57.9 my_nb 17079 63.27 el_sw 12682 65.42
fi_uk 60369 67.11 fr_mr 39613 59.97 fr-ca_lt 28019 66.5 ar_az 22828 54.89 bs_de 19763 69.36 bn_cs 17031 71.02 gl_mk 12643 68.66
gl_ja 60292 61.48 et_zh-tw 39470 59.6 de_mr 27919 60.14 eo_it 22807 65.94 ta_uk 19755 60.35 kk_th 17025 58.53 en_uz 12605 68.42
cs_sl 60257 64.26 et_it 39396 66.73 gu_ko 27897 65.42 fa_gu 22775 68.07 bn_nl 19743 71.43 eo_sr 16994 64.07 et_hy 12568 68.97
gl_he 60181 68.29 ar_mn 39327 53.35 fa_ur 27841 64.79 bs_es 22773 71.13 es_eu 19707 63.62 pt-br_sw 16991 68.23 it_uz 12541 58.68
nb_pt-br 59946 70.65 et_ko 39184 60.15 ar_gu 27822 65.92 gu_tr 22694 67.19 he_kk 19685 60.69 bg_zh 16956 55.2 ja_uz 12530 59.06
gl_vi 59715 66.1 sk_sl 39176 64.98 ta_zh-tw 27729 58.35 eo_tr 22681 57.26 ar_sw 19635 61.62 mr_sk 16946 64.61 lt_ur 12494 64.75
hy_sr 59393 71.74 et_ru 39081 66.39 da_fr-ca 27718 70.59 ja_zh 22668 52.78 id_ta 19626 64.77 az_hr 16927 61.06 sq_ta 12451 59.64
ja_nb 59370 61.38 et_tr 39034 57.95 gl_sv 27669 69.53 bs_hu 22665 66.94 eu_ko 19608 60.61 sw_tr 16892 59.86 fr-ca_kmr 12444 64.61
cs_my 59125 60.86 fa_mr 38967 59.34 ta_tr 27550 61.7 de_ta 22650 63.8 bn_de 19583 67.93 ja_sw 16827 60.78 uz_vi 12421 54.52
da_sq 59094 65.53 et_zh-cn 38912 57.51 es_ta 27516 63.76 hu_ms 22587 62.27 de_eo 19563 63.3 gu_id 16791 69.13 bs_sk 12421 68.98
hy_th 59014 67.56 my_sk 38860 61.54 ca_mk 27510 68.37 ja_ms 22565 59.52 ca_gl 19537 65.35 cs_zh 16676 55.99 hi_ta 12404 65.6
gl_it 58966 69.96 ka_ro 38839 65.8 hr_mn 27478 56.43 it_zh 22539 58.55 az_en 19496 67.92 mr_pt 16585 64.5 pl_sw 12365 63.34
da_lt 58519 67.93 he_mr 38829 59.79 ka_sv 27471 69.57 ms_pl 22476 64.07 eo_hu 19488 63.85 az_th 16562 57.18 gl_sl 12339 66.12
ckb_hu 58034 64.8 mr_tr 38771 59.25 et_id 27424 63.71 ar_eo 22475 59.18 eu_fr 19470 63.44 ka_pt 16527 68.13 fr_uz 12310 56.14
en_fr-ca 57838 77.03 et_fr 38519 64.46 ko_ur 27368 63.68 gu_pt-br 22474 70.83 sw_zh-tw 19459 62.71 mk_my 16510 59.47 az_sk 12257 61.21
en_nb 57805 79.09 he_mn 38466 55.85 ur_zh-tw 27255 63.45 nl_ta 22441 65.43 fi_ka 19432 63.92 fr-ca_lv 16509 64.06 mr_sq 12256 58.03
mk_th 57666 63.23 ka_sr 38408 68.84 mr_uk 27138 61.46 bs_fr 22433 71.93 eu_zh-cn 19430 54.12 da_et 16428 67.72 pt_ur 12243 71.63
fr-ca_ro 57611 67.98 es_mr 38366 60.75 ca_lv 27123 61.62 gu_hu 22393 67.16 id_ur 19426 71.55 gl_hi 16407 66.6 et_lv 12218 60.78
nb_pl 57576 68.01 lv_sv 38141 66.86 tr_ur 27086 63.83 eo_zh-tw 22319 60.18 eu_tr 19418 57.48 th_zh 16333 56.85 fr-ca_my 12206 61.04
gl_hu 57234 66.87 mn_ru 38121 54.51 he_ta 27010 59.85 ms_pt-br 22286 65.62 az_ro 19418 56.86 sw_zh-cn 16325 60.37 be_cs 12195 66.89
hy_uk 57028 68.8 ko_mn 38082 57.55 mr_th 26991 56.83 bn_ko 22276 64.43 eo_pt-br 19397 66.27 eo_uk 16320 64.88 gu_hr 12183 72.61
fi_id 56914 64.69 et_ja 38026 57.93 ta_zh-cn 26931 56.98 az_es 22258 58.49 gu_he 19393 66.39 kk_nl 16284 64.74 bn_sv 12169 70.14
sl_th 56724 66.49 lt_sl 37983 60.39 ta_vi 26903 60.73 bn_he 22257 65.49 bn_hr 19389 71.84 ca_ka 16175 67.12 ar_ml 11966 60.09
hu_nb 56600 67.12 bg_et 37961 65.81 ka_sk 26776 67.12 az_fa 22254 59.82 lv_sl 19365 62.39 kk_pt-br 16119 63.87 sk_ur 11937 68.43
gl_pt-br 56578 67.96 ka_uk 37846 66.3 da_my 26764 64.27 ka_sq 22240 62.36 eu_ja 19355 58.77 uk_zh 16116 54.37 sv_zh 11928 59.52
lv_th 56527 65.73 et_pt-br 37667 65.18 fa_ta 26706 61.89 ckb_cs 22163 68.76 ckb_sv 19351 67.85 hi_sl 16112 70.5 bn_hi 11924 66.91
pt_sq 56464 66.31 es_et 37614 64.9 el_mn 26676 54.79 uk_ur 22142 66.13 az_el 19316 58.55 hy_nb 16068 71.38 sv_ta 11922 66.31
fa_nb 56419 69.16 et_pl 37564 64.03 bg_mn 26670 51.37 kmr_sl 22102 61.45 hr_ta 19224 66.11 eo_id 16054 65.23 eo_lt 11836 59.67
el_fr-ca 56408 71.27 hi_my 37548 63.11 da_hi 26669 69.97 az_tr 22094 63.69 gu_ja 19218 65.61 kmr_mr 16046 51.64 sw_uk 11826 66.39
cs_hi 56176 68.12 et_he 37462 62.98 en_ur 26655 76.77 eo_he 22093 65.22 be_zh-tw 19178 56.75 az_id 16033 59.39 bg_gu 11818 68.23
bg_nb 55914 71.8 mn_zh-tw 37406 53.22 ko_ta 26604 61.38 bn_fr 22070 69.88 fa_sw 19173 66.28 be_pt-br 15982 65.6 ms_pt 11789 66.75
cs_fr-ca 55450 68.81 hy_sv 37366 72.18 gu_ru 26514 67.08 bn_zh-cn 22057 60.76 ko_sw 19124 61.84 en_eu 15948 73.52 hu_uz 11755 56.34
fr-ca_hr 55429 72.75 cs_ka 37300 64.56 fr_ur 26319 69.22 bs_en 21998 79.63 ms_sr 19102 66.89 eu_hr 15942 63.02 be_id 11737 64.08
id_mk 55413 70.51 mr_zh-cn 37210 56.63 fr_ta 26235 62.64 fa_zh 21985 55.54 eo_fa 19099 63.13 eo_th 15885 62.9 ml_zh-tw 11695 59.21
de_nb 55166 68.54 ja_mr 36933 57.31 nb_sq 26220 65.87 bs_vi 21947 69.69 kk_ko 19078 64.2 hr_zh 15833 58.34 pt-br_uz 11625 57.15
id_sl 54741 66.25 ja_mn 36928 58.89 ar_ms 26168 64.18 gl_sk 21942 67.17 be_it 19061 66.37 hy_my 15825 62.51 ka_nb 11598 67.74
nb_nl 54342 69.03 en_mn 36906 66.06 fr_gu 26152 69.1 eo_ru 21921 64.08 my_sl 19049 61.26 be_nl 15783 66.17 fr-ca_gl 11568 64.39
hi_pt 54241 71.39 et_vi 36883 63.49 it_ta 26123 65.49 az_ko 21897 62.53 nl_zh 19021 59.61 lt_mr 15776 57.9 ckb_sk 11560 66.77
ca_sk 54232 66.05 mn_vi 36734 54.09 es_ur 26057 70.59 az_ja 21888 58.95 et_sv 19009 66.7 hi_nb 15738 72.79 bs_sq 11559 62.87
lt_sq 54156 56.71 et_hu 36657 61.03 fr-ca_pt 25925 68.34 az_fr 21880 58.92 eo_hr 18989 65.3 ca_nb 15731 66.5 th_uz 11534 56.71
lt_pt 53925 63.74 it_mr 36650 62.65 et_th 25852 64.12 az_he 21849 58.59 be_ru 18988 64.69 ca_et 15707 61.66 ml_tr 11495 61.02
ca_sv 52925 66.43 cs_gl 36604 62.84 cs_mn 25800 54.98 sr_ur 21795 70.89 bn_pl 18970 68.23 be_pl 15631 65.02 bg_sw 11463 63.24
fr-ca_uk 52898 66.73 sl_sq 36579 58.68 et_uk 25797 68.09 eo_es 21737 64.21 cs_ur 18914 70.37 az_sr 15629 60.31 eo_sk 11450 65.18
gl_pl 51375 66.6 mn_tr 36547 57.31 gu_zh-tw 25733 65.37 bs_nl 21720 70.36 ru_sw 18901 63.09 kk_sr 15552 63.45 ka_sl 11429 65.76
ckb_pl 51099 66.74 fa_mn 36461 55.77 it_ur 25707 71.71 az_vi 21692 56.94 kk_tr 18899 61.57 be_th 15549 61.16 lv_mn 11406 53.06
el_nb 51088 72.82 fr-ca_sq 36388 61.53 ar_zh 25696 51.97 bn_fa 21691 64.71 fi_nb 18779 69.15 bg_kk 15478 57.82 hy_ka 11394 65.98
ka_ru 51082 65.91 et_nl 36312 66.24 hu_ur 25645 68.3 gl_lt 21685 63.15 kk_zh-tw 18769 58.06 eu_uk 15408 61.18 ca_ckb 11366 66.36
gl_nl 50958 69.47 hu_mr 36226 59.99 he_ur 25522 67.7 pl_ur 21638 67.81 eu_nl 18749 64.21 hu_sw 15407 64.91 ko_ml 11353 59.43
fr-ca_id 50569 68.86 it_mn 36199 58.48 hy_pt 25483 71.41 hy_lv 21634 66.56 hy_sl 18727 69.37 az_uk 15389 60.31 ms_sk 11302 65.45
ckb_pt 50551 68.2 de_et 36069 64.42 ca_kmr 25446 65.15 id_ms 21607 63.96 eu_he 18677 63.2 de_kk 15385 62.4 de_sw 11236 65.05
fr-ca_sr 50486 70.32 kmr_lt 36027 61.76 ca_fi 25368 64.53 az_zh-cn 21597 54.2 bs_hr 18627 70.46 mn_my 15331 57.34 eu_sq 11226 53.6
ar_ka 50441 63.95 hi_sk 36021 72.27 en_ta 25349 70.7 el_ta 21541 61.98 az_cs 18597 61.77 et_lt 15323 59.79 el_uz 11176 57.28
da_pt 50372 70.39 ka_th 36008 64.53 ar_bs 25241 69.34 kmr_mk 21520 53.46 lv_my 18587 61.26 it_sw 15277 68.68 az_sv 11150 60.55
id_lv 50223 63.35 da_mk 35959 71.34 es_gu 25191 70.82 et_sk 21501 65.13 ms_uk 18575 67.31 be_uk 15276 64.74 ml_ru 11103 60.45
ka_zh-tw 49764 61.37 mr_vi 35559 57.6 id_mn 25142 58.71 bs_pl 21487 66.6 en_kk 18564 69.69 be_sr 15265 65.62 ca_mn 11093 54.99
ka_ko 49635 60.09 es_mn 35453 56.74 ja_ur 25094 60.9 az_bg 21478 58.18 hi_mr 18551 60.63 id_zh 15236 58.39 da_mr 11076 62.53
it_ka 49548 68.79 fr_mn 35254 56.27 pt-br_ur 25040 71.93 az_hu 21458 58.93 ms_ro 18534 65.96 gl_kmr 15186 66.36 bn_sk 11009 69.75
nb_ro 48567 69.7 fi_lt 35186 60.71 ca_sl 25001 65.22 ro_ta 21406 61.09 en_eo 18522 73.46 gu_nl 15149 71.25 sq_ur 10964 66.61
ka_tr 48302 63.42 hu_mn 34911 55.17 hu_ta 24940 62.63 ka_lt 21391 63.7 fa_kk 18479 60.85 bs_id 15067 69.35 ckb_gl 10956 63.79
ka_vi 48232 65.67 lt_my 34686 59.18 ca_hi 24937 66.57 az_de 21378 57.65 cs_eo 18457 60.29 fr-ca_mk 15030 68.18 hi_ka 10862 67.74
gl_ro 48142 67.78 mk_sq 34644 64.06 cs_mr 24894 61.27 bn_ja 21359 61.89 fr-ca_hy 18456 70.27 el_gu 15024 68.07 bn_lt 10858 65.42
ca_sq 47924 61.82 ckb_de 34599 62.32 bs_zh-tw 24858 65.05 th_ur 21187 64.18 cs_ms 18400 61.98 sw_vi 14972 65.02 es_ml 10854 64.13
fa_ka 47321 64.32 lt_lv 34391 60.49 ms_ru 24832 65.3 eo_ko 21185 59.69 sr_zh 18377 56.76 eu_ro 14948 60.69 lt_ta 10826 61.19
es_ka 47230 66.3 mn_zh-cn 34171 51.71 gu_zh-cn 24827 62.46 eo_ja 21169 57.97 eu_pl 18330 60.94 ms_sq 14928 59.48 sr_uz 10813 56.37
ja_ka 47197 60.48 fr-ca_sk 34167 69.6 bs_ru 24748 69.32 bg_ta 21147 61.95 fi_hy 18284 67.29 be_bg 14909 62.06 gl_hy 10803 70.41
fr_ka 47174 67.59 da_sl 34079 69.35 ur_vi 24688 65.15 pt-br_zh 21088 56.81 fr_sw 18270 64.8 eu_th 14889 60.24 it_ml 10769 64.81
id_nb 47086 70.81 kmr_sq 33980 55.89 ms_zh-tw 24671 62.67 fi_kmr 21069 64.69 be_en 18247 71.88 cs_kk 14847 61.78 ja_ml 10749 61.16
kmr_pt 46732 67.43 mr_pt-br 33544 64.28 da_nb 24664 70.06 eo_zh-cn 21045 57.76 be_ko 18241 58.19 he_sw 14819 67 ka_kmr 10723 64.74
hr_nb 46727 69.37 ckb_nl 33495 67.63 ur_zh-cn 24643 61.36 en_ms 21031 75.62 kk_zh-cn 18211 56.42 be_ro 14802 64.1 hr_uz 10721 57.05
ckb_my 46678 54.56 mr_ro 33260 62.72 ms_vi 24611 67.06 az_pl 21031 60.58 gl_my 18195 63.19 be_el 14749 65.89 bn_da 10711 70.52
ka_nl 46444 67.78 hy_lt 33201 66.4 he_ms 24611 64.01 fi_hi 20962 67.36 eu_vi 18178 60.24 kk_ro 14714 60.82 uk_uz 10709 56.22
he_ka 46357 61.79 mn_th 33067 55.37 ckb_hi 24607 57.75 pl_zh 20953 55.04 mr_my 18160 57.97 mn_pt 14680 57.52 id_uz 10667 56.91
hu_ka 46319 64.54 mn_pl 33052 58.32 ms_zh-cn 24569 60.58 bg_bs 20951 72.58 el_zh 18158 55.9 fi_fr-ca 14570 67.65 fa_ml 10626 62.81
my_sv 46202 60.36 lt_mk 32864 64.01 pl_ta 24545 60.95 el_ms 20867 66.26 ca_fr-ca 18137 65.97 gu_sr 14540 72.69 ca_mr 10614 59.34
ka_zh-cn 46162 59.64 da_lv 32799 65.36 da_ka 24416 66.89 eo_nl 20844 67.14 be_he 18084 62.25 id_kk 14453 65.57 bn_sq 10588 64.4
ckb_sr 45921 65.44 lv_sq 32747 55.62 ar_bn 24390 64.86 de_ms 20832 64.34 de_zh 18083 55.45 kmr_nb 14434 65.87 nl_uz 10578 60.57
hy_sk 45752 71.1 da_hy 32704 71.13 pt-br_ta 24343 64.22 ta_th 20802 60.6 bs_sr 18083 70.81 eu_sr 14417 62.04 bs_pt 10545 70
kmr_my 45744 54.78 el_et 32632 66.25 fa_ms 24312 63.97 de_ur 20795 68.49 sr_ta 18060 64.65 ar_uz 14404 54.29 eu_sv 10534 63.75
en_mr 45730 66.97 fi_sq 32527 59.1 vi_zh 24297 56.2 bg_ur 20787 68.36 be_tr 18051 59.91 kmr_mn 14368 52.06 en_ml 10519 71.93
cs_nb 45384 67.23 cs_et 32279 63.18 bs_ko 24288 64.4 ms_nl 20732 67.56 eu_pt-br 18047 64.79 ka_mk 14346 65.87 fr_ml 10466 64.62
gl_sr 45311 70.74 gl_pt 32122 68.61 zh_zh-tw 24269 56.7 eo_pl 20721 63.37 bn_th 17989 64.78 kk_uk 14321 59.77 az_da 10459 59.48
ka_pl 45245 65.32 ckb_uk 31970 64.33 bg_ckb 24250 63.86 fr-ca_sl 20670 67.43 be_ja 17965 56.77 lv_nb 14298 65.82 de_uz 10448 56.76
fr-ca_th 45085 65.45 et_fa 31890 64 ms_tr 24249 61.09 ro_zh 20663 54.99 bn_ro 17958 68.68 el_kk 14280 61.06 eo_fi 10448 61.71
fi_sk 45056 65.28 da_kmr 31836 68.38 ja_ta 24230 61.23 hy_kmr 20645 63.61 es_kk 17949 61.94 et_sq 14174 56.76 et_fi 10435 64.29
lv_sk 44698 63.42 mr_pl 31823 64.55 bs_he 24222 69.51 ar_eu 20624 57.98 be_hu 17929 63.11 el_eu 14136 60.19 sk_zh 10426 56.25
gl_hr 44676 70.75 hi_lt 31813 69.49 es_ms 24177 64.65 bn_en 20597 78.87 ja_kk 17915 62.1 fr-ca_nb 14104 69.95 et_mk 10424 63.95
kmr_sv 44578 66.39 et_ro 31565 61.6 zh_zh-cn 24153 55.42 hu_zh 20593 55.7 hr_ms 17906 64.92 nb_sl 13969 68.74 hr_sw 10394 65.97
el_gl 44371 69.51 en_et 31491 76.05 kmr_lv 24114 61.91 bg_eo 20521 62.6 bs_th 17884 69.34 nl_sw 13955 69.69 kmr_ur 10387 53.82
de_gl 44351 66.14 ckb_hr 31415 67.3 ko_ms 24109 57.63 gl_sq 20505 62.75 fr_kk 17851 60.88 gu_pl 13947 67.87 pt_ta 10386 66.09
ar_mr 44345 55.99 nb_sk 31331 70.09 fr_ms 24041 65.14 bs_el 20496 73.49 gu_vi 17846 68.02 cs_eu 13817 58.17 lt_ms 10355 57.6
ka_pt-br 44224 67.69 mn_pt-br 31262 57.8 fi_mk 23944 66.85 bn_pt-br 20484 70.62 gu_ro 17813 67.82 uz_zh-tw 13795 55.87 pl_uz 10333 56.04
bg_gl 44036 69.44 my_sq 30974 55.95 fi_sl 23939 64.12 bg_ms 20445 66.7 bg_eu 17800 58.13 hr_kk 13782 63.83 kk_sk 10265 63.56
fi_sv 43981 68.78 et_hr 30957 66.32 ko_zh 23783 54.57 ar_be 20377 61.87 bs_uk 17755 69.39 pt_zh 13733 57.49 be_sv 10244 66.78
da_fi 43889 68.79 hi_sq 30884 65.66 fi_lv 23775 59.66 hr_ur 20299 71.28 de_eu 17754 59.78 fi_gl 13725 62.88 ml_pt-br 10239 63.11
nb_sr 43859 72.41 mn_nl 30833 59.13 bn_ru 23741 66.04 eu_zh-tw 20289 57.44 bs_cs 17731 68.23 uz_zh-cn 13676 52.8 et_fr-ca 10121 65.85
ckb_th 43835 55.83 nb_sv 30790 71.85 en_zh 23704 65.09 el_ur 20251 68.97 el_eo 17717 63.68 gl_lv 13635 63.06 ckb_mr 10118 52.14
bg_ka 43727 65.54 lv_pt 30777 63.73 bn_zh-tw 23673 63.3 eo_fr 20250 65.53 be_es 17716 64.75 be_hr 13614 66.99 pt_sw 10103 67.98
mk_sk 43549 70.95 mr_nl 30714 64.87 lt_nb 23606 65.41 az_pt-br 20229 59.09 bn_id 17662 70.42 sr_sw 13592 65.34 ml_zh-cn 10073 56.24
en_ka 43546 74.21 mn_uk 30417 54.76 tr_zh 23581 56.13 ro_ur 20219 70.49 es_sw 17661 67.03 ckb_lt 13526 62.11 eu_sk 10040 60.75
ca_lt 43456 63.16 fi_pt 30314 66.48 bs_ja 23501 62 eu_it 20180 65.92 mk_sl 17659 67.73 mn_sq 13463 51.2 da_ta 10000 66.58
kmr_sk 42999 65 fr-ca_sv 30201 70.31 ca_my 23454 60.34 az_nl 20172 61.74 eu_hu 17622 60.16 ru_uz 13402 56.39 hi_mn $99$ 59.44
gl_uk 42801 69.01 id_ka 30092 66.66 ru_zh 23421 53.06 eu_ru 20153 62.9 eu_fa 17593 62.06 tr_uz 13382 57.17 ca_ta 9983 62.74
gl_th 42597 66.15 mk_pt 29821 72.6 it_ms 23406 65.38 da_gl 20150 66.6 hu_kk 17567 61.42 de_gu 13380 68.87 bg_uz 9961 55.67
hi_sv 42572 72.89 en_gu 29681 79.07 bn_es 23368 69.07 bn_hu 20096 67.33 eo_ro 17554 63.54 eu_id 13361 61.51 gu_hi 9942 71.87
ckb_el 42146 64.63 el_mr 29645 60.26 bn_tr 23346 64.51 ar_kk 20094 57.68 hi_hy 17551 71.71 sv_ur 13327 71.6 az_lt $98$ 55.99
de_ka 42146 65.67 mn_sr 29516 56.99 he_zh 23332 52.33 cs_ta 20071 62.89 be_vi 17466 60.74 ms_sv 13297 67.43 da_eu 9856 62.87
gl_id 42075 67.71 mr_sr 29287 63.78 bs_zh-cn 23275 62.43 hy_mk 20052 68.43 mk_nb 17459 71.62 gu_pt 13197 72.2 kk_my 9850 60.32
mk_sv 41537 71.95 et_sr 29282 65.05 bn_it 23263 71.99 bs_ro 20025 69.38 bn_sr 17407 70.48 ro_sw 13165 65.85 cs_uz 9783 55.03
Table 13: COMET-QE score and the count of sentences for all 4,765 language pairs in TED2025 (part II).
Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE
gl_nb 9763 69.76 it_ne 7276 68.96 ne_uk 5902 67.79 ca_sw 4657 64.38 kn_zh-tw $39$ 57.23 en_kn 2984 69.3 ka_uz 2385 57.03
sw_th 9717 64.26 fil_tr 7235 55.82 eo_lv 5847 60.91 ar_sh 4655 68.13 de_kn 3937 62.02 bn_ckb 2975 64.63 fi_ne 2380 66.46
lv_mr 9713 60.05 fil_ja 7210 57.82 id_ml 5817 63.08 az_my 4643 62.26 cs_srp 3933 68.93 cs_si 2972 63.69 ja_tl 2370 58.68
ckb_sq 9677 56.9 eu_mk 7177 59.87 fil_id 5816 61.9 ckb_gu 4633 62.64 pt_srp 3924 71.02 cs_kn 2971 58.32 de_tl 2361 61.46
kk_pt 9660 64.35 te_th 7167 60.45 ne_th 5800 62.89 is_zh-tw 4615 58 kn_tr 3923 61.92 mn_uz 2970 54.7 pt-br_tl 2361 62.52
fi_mr 9659 61.05 et_kmr 7154 59.83 eu_pt 5798 64.97 sr_srp 4608 71.38 is_pl 3913 61.64 bs_nb 2969 70.56 el_kn 2357 58.97
ar_te 9648 57.75 sq_uz 7149 49.5 az_kmr 5783 54.73 hu_is 4598 59.21 bg_kn 3904 60.9 lt_ml 2964 57.47 eu_mn 2354 52.06
az_sq 9604 53.53 gu_kmr 7124 63.73 ru_srp 5779 69.01 gu_sk 4595 69.04 hr_kn 3902 63.07 et_kk 2935 59.62 en_nn 2353 78.32
be_pt 9569 65.05 es_fil 7116 61.7 ne_pl 5764 71.22 hy_uz 4586 60.16 si_th 3885 66.56 et_ms 2926 61.36 gu_mk 2333 68.46
da_eo 9527 64.91 fil_ru 7116 62.01 si_vi 5758 65.81 is_vi 4581 62.97 kn_ko 3872 59.7 fil_hi 2919 62.84 arq_vi 2325 49.95
mn_sl 9527 55.84 eu_fi 7114 59.51 gu_sv 5750 71.19 gl_ms 4574 64.76 gl_gu 3870 70.56 bs_ta 2914 66.79 nl_tl 2316 65.27
fi_mn 9525 55.82 ckb_fi 7103 64.57 be_lv 5737 61.45 is_tr 4574 59.29 nl_si 3865 68.78 nn_zh-cn 2911 61.28 lv_ml 2316 57.06
hu_ml 9498 61.66 fil_pt-br 7103 63.93 el_fil 5728 62.71 sq_sw 4573 59.73 cs_so 3864 45.14 arq_en 2911 66.15 ja_nn 2312 61.3
te_tr $94$ 61.52 ckb_mn 7092 51.96 hi_te 5722 63.4 kk_mr 4569 57.44 ms_ur 3859 69.28 eo_mr 2902 57.31 et_fil 2310 54.28
be_sk 9480 66.02 az_pt 7046 60.25 my_uz 5706 59.45 is_it 4563 65.91 he_si 3858 62.48 kn_th 2883 63.67 ca_is 2309 61.73
et_sl 9456 63.91 fil_vi 7044 61.25 fr-ca_ms 5705 64.66 ckb_zh 4560 53.24 da_ne 3856 68.24 ml_sl 2883 62.81 az_kk 2307 60.95
hi_ur 9437 67.83 hi_sw 7040 63.17 bn_hy 5674 68.58 fil_lt 4539 56.82 be_ka 3848 60.36 it_nn 2875 68.03 hi_kn 2298 65.96
ca_ur 9410 66.46 nl_te 7032 65.38 pl_srp 5660 66.07 ne_ro 4538 69.61 eu_gl 3829 61.49 is_sk 2859 64.42 eu_ta 2291 65.15
de_ml 9288 63.38 eo_fr-ca 7027 64.87 bn_ka 5659 65.63 it_so 4533 50.67 sw_ur 3826 63.29 nn_tr 2858 63.63 is_uk 2289 64.13
my_ur 9268 60.22 fil_fr 7023 64.04 cs_ne 5651 67.02 he_so 4533 49 mk_zh 3817 54.45 arq_pt-br 2851 56.19 arq_id 2288 55.13
ml_nl 9258 63.7 fil_hu 7022 61.23 pt-br_si 5640 68.48 ar_so 4529 46.84 en_so 3779 62.39 hu_nn 2843 64.9 nn_sq 2272 64.27
hy_mn 9234 56.11 ml_uk 7022 61.36 hr_ne 5637 68.08 is_ko 4515 60.02 fa_kn 3771 63.31 bn_te 2832 63.3 mk_ml 2272 59.35
ca_eu 9229 60.19 ne_vi 7020 63.39 ka_mr 5619 59.93 srp_th 4513 69.8 el_is 3765 63.58 it_tl 2830 63.47 nn_th 2265 69.8
en_te 9221 71.9 be_sl 7008 62.64 eo_ka 5618 62.89 pl_sh 4502 66.51 sl_uz 3758 55.19 nn_ru 2825 66.42 arq_de 2259 47.72
id_sw 9203 67.71 he_ne 6991 65.35 ckb_lv 5614 63.74 ja_so 4497 50.18 el_sh 3756 73.69 bs_mn 2823 57.23 fr-ca_ml 2256 64.93
fa_te 9189 61.92 fil_he 6977 63.19 fil_uk 5607 62.51 id_si 4487 68 az_mn 3753 53.99 el_nn 2818 72.06 so_sv 2249 51.43
ru_te 9168 58.25 bs_hi 6977 69.55 be_ca 5602 60.83 gl_uz 4486 56.56 az_bn 3752 62.27 arq_hu 2818 55.33 kmr_ml 2246 59.69
ko_te 9167 59.29 gl_ka 6977 66.59 fa_si 5600 65.24 da_ml 4483 64.64 mk_uz 3745 55.09 be_ur 2816 61.83 gl_sw 2239 63.48
kk_sv 9116 64.09 et_my 6974 60.52 bn_lv 5598 67.37 et_mn 4473 52.69 so_uk $37$ 52.07 en_is 2804 74.55 arq_cs 2239 49.84
bn_ca 9107 66.4 az_lv 6924 57.09 ms_sl 5598 62.21 fa_sh 4466 70.29 ca_ne 3729 65.6 sh_sq 2792 60.86 az_fil 2229 52.96
ca_ms 9106 61.66 hu_ne 6922 65.96 ml_th 5576 61.91 ml_sv 4465 63.82 ca_gu 3719 69.72 arq_fr 2779 53.37 fa_ps 2220 59.02
fi_ta 9066 62.97 ne_tr 6915 64.86 en_fil 5565 73.06 gl_ur 4454 67.26 sh_th 3712 66.5 fr_nn 2776 65.69 ar_km 2199 62.73
ro_uz 9042 53.83 ne_ru 6914 68.6 mr_ur 5557 64.18 eu_nb 4447 63.28 az_ta 3708 64.66 arq_nl 2773 56.07 gu_mn 2197 60.96
bs_lt 9039 63.82 bn_nb 6908 72 pt-br_srp 5536 71.95 ko_so 4445 49.66 et_ta 3708 62.61 nn_zh-tw 2772 62.36 tl_tr 2196 58.2
hi_ms 9023 69.93 de_fil 6891 61.61 hy_kk 5524 61.23 so_zh-cn 4433 46.87 ckb_si 3707 63.92 eo_fil 2759 57.68 fil_nb 2193 64.34
te_zh-tw 9022 58.65 az_sl 6891 59.57 eo_kmr 5504 58.04 ckb_kk 4426 53.5 kn_pl 3701 59.52 eo_ta 2752 64.62 en_ps 2182 71.83
ml_vi 9013 63.18 be_da 6885 64.53 bs_lv 5496 64.56 es_so 4415 47.38 so_sr 3701 51.25 is_sq 2745 57.39 hy_is 2181 63.05
ml_pl 9006 60.12 fa_ne 6885 63.45 si_zh-tw 5496 63.22 fil_sk 4414 61.87 fil_hy 3699 66.13 arq_it 2743 54.76 fr-ca_ne 2165 64.12
et_pt 8999 65.64 ar_si 6880 64.73 az_ka 5461 58.72 fr_so 4403 47.75 ms_ta 3693 64.88 eo_et 2737 59.51 bs_eo 2163 60.06
ml_sr 8940 64.36 fa_fil 6841 58.77 bs_sl 5456 67.93 bn_fr-ca 4388 68.25 he_kn 3671 58.27 sq_srp 2735 62.35 si_sv 2158 68.37
cs_sw 8917 65.34 ms_nb 6836 66.83 az_nb 5423 60.32 gl_kk 4386 63.42 is_th 3664 62.56 be_et 2732 62.08 lt_so 2151 43.18
gu_th 8879 66.36 bn_my 6812 63.28 eu_sl 5411 59.44 ckb_ms 4381 65.43 kn_nl 3644 63.03 be_fr-ca 2730 63.41 da_te 2150 61.59
my_zh 8849 54.59 es_ne 6810 67.61 kk_nb 5397 64.34 be_hy 4381 64.09 is_ja 3641 56.11 fil_pt 2719 63.56 mn_sw 2146 54.32
cs_gu 8826 67.74 kk_kmr 6798 54.37 sk_sw 5377 64.5 is_pt-br 4381 64.74 kk_ms 3639 62.53 hr_nn 2719 65.51 km_ko 2140 63.12
it_te 8776 64.7 da_kk 6747 63.13 bn_ur 5301 66.49 ja_si 4375 61.36 ka_kk 3639 60.27 bg_nn 2713 69.08 nn_pt-br 2134 68.4
eu_lt 8759 56.03 fil_ko 6726 58.79 eu_fr-ca 5299 62.61 de_is 4370 63.98 fa_is $36$ 61.57 cs_nn 2700 63.58 is_lv 2132 59.08
bs_da 8743 69.14 en_ne 6725 73.58 fr_si 5297 67.83 ru_so 4344 46.47 hr_si 3632 67.23 ta_te 2684 58.71 ja_ps 2125 61.54
da_ur 8722 70.01 it_srp 6719 72.56 cs_te 5297 60.27 ru_sh 4328 70.13 sh_sr 3628 69.18 tl_zh-cn 2672 56.19 ko_tl 2119 58.26
be_my 8705 56.53 lv_ur 6708 66.35 fr-ca_mr 5293 59.47 cs_is 4323 59.35 ka_ms 3587 62.91 bs_ka 2658 67.08 ps_zh-tw 2114 60.79
eo_sq 8658 56.17 ne_zh-tw 6705 64.24 hi_kk 5291 62.69 he_sh 4311 70.64 az_et 3579 55.05 ka_zh 2654 54.36 arq_hr 2109 50.96
ca_eo 8657 60.68 hy_mr 6690 61.47 mr_ta 5271 61.27 hy_ur 4303 69.75 be_mk 3546 60.64 id_kn 2652 61.07 ps_ru 2104 66.67
mk_mn 8647 50.65 fil_zh-cn 6689 54.67 kk_sl 5270 61.53 be_mn 4303 49.49 ml_sq 3546 55.79 arq_el 2643 56.61 da_nn 2103 68.84
es_te 8602 63.29 te_uk 6686 60.68 el_te 5261 60.41 ckb_uz 4297 52.92 si_uk 3542 65.88 fi_te 2641 60.67 ar_ps 2102 56.35
be_lt 8545 61.52 ko_ne 6672 62.13 be_hi 5256 62.45 ne_sk 4295 71.93 hu_kn 3534 62.1 id_tl 2636 61.04 lt_srp 2102 63.89
fr_te 8526 62.58 da_zh $66$ 57.72 eu_lv 5248 56.42 ka_ur 4279 67.27 fr-ca_uz 3527 56.58 es_tl 2623 62.52 km_tr 2097 62.26
da_ms 8506 65.29 ar_srp 6661 70.41 fr-ca_mn 5225 55.77 hu_si 4269 62.21 mr_ms 3525 62.85 bn_ms 2618 69.58 ro_tl 2093 59.58
he_te 8478 58.99 es_srp 6647 72.91 az_mk 5213 57.22 lt_te 4267 58.13 be_nb 3512 64.9 sw_ta 2613 60.2 en_km 2090 78.05
ms_my 8476 62.03 nb_ta 6624 66.56 bn_ta 5198 64.47 pl_so 4265 46.23 bg_si 3506 66.38 arq_ja 2609 51.17 bn_uz 2083 61.6
pl_te 8468 60.89 lt_uz 6622 51.42 it_si 5187 67.98 is_ru 4257 62.45 sv_te 3503 65.14 arq_zh-cn 2608 47.98 kn_ur 2083 63.91
ja_te 8463 60.84 da_uz 6553 58.93 ro_srp 5187 69.23 so_zh-tw 4251 50.09 fa_so 3501 49.75 bn_eo 2608 65.44 bn_kn 2083 66.38
kk_lt 8461 58.21 cs_fil 6533 59.86 ml_sk 5157 61.95 fil_fr-ca 4248 63.86 eu_kmr 3491 56.93 arq_uk 2605 53.34 kn_sw 2083 63.8
bg_ml 8458 61.88 fi_ms 6505 60.77 sl_ur 5150 67.65 ar_is 4239 60.34 et_eu 3455 52.62 fa_nn 2604 66.49 hr_tl 2083 61.47
te_vi 8370 60.57 ca_uz 6505 53.81 lt_sw 5150 59.09 az_eo 4239 54.42 gu_lv 3450 66.86 fil_lv 2602 56.91 eo_zh 2079 56.91
el_ml 8362 61.41 en_srp 6496 80.08 pt_uz 5144 57.56 fil_sv 4235 65.45 de_si 3450 66.87 lv_sw 2596 61.35 bo_vi 2076 41.92
be_sq 8345 58.57 sl_ta 6495 62.93 fil_sq 5140 54.75 hu_so 4231 47.68 et_mr 3436 58.72 ca_ml 2594 63.44 bo_fr 2076 40.91
bs_ca 8334 66.61 ja_ne 6447 63.47 da_gu 5140 69.81 ja_sh 4230 62.09 sk_te 3433 60.02 id_nn 2587 66.97 bg_bo 2076 36.78
kmr_ta 8324 59.68 et_gl 6444 63.45 nb_zh 5137 59.92 ckb_ta 4226 58.88 kn_ro 3427 58.55 he_tl 2585 62.14 bo_es 2076 41.23
he_ml 8321 59.41 hi_zh 6441 60.82 da_sw 5127 69.76 ka_ta 4218 63.04 be_mr 3424 57.09 nn_sk 2583 67.37 ar_bo 2076 37.68
bn_pt 8306 71.45 lv_ms 6430 60.02 id_srp 5126 70.81 bn_mn 4218 58.6 id_so 3422 48.83 he_nn 2583 66.91 bo_zh-tw 2076 41.03
sq_zh 8291 54.25 fi_ur 6430 66.86 ko_si 5125 62.93 az_fr-ca 4215 58.18 hi_ml 3401 65.66 my_srp 2574 64.39 bo_ja 2076 47.23
hy_ms 8274 66.03 fil_nl 6425 65.73 ml_pt 5110 64.37 ckb_hy 4202 63.83 so_th 3393 53.51 ka_ne 2567 66.05 bo_it 2076 42.41
ro_te 8244 61.92 ne_nl 6403 71.42 ta_ur 5071 61.9 bn_gl 4202 67.8 az_ur 3392 61.85 hu_tl 2556 60 bo_zh-cn 2076 38.95
kmr_ms 8237 62.41 es_si 6400 68.5 fil_ro 5065 61.59 bg_so 4192 45.88 et_uz 3388 52.36 az_zh 2549 53.43 bo_he 2076 41.33
mn_nb 8203 56.35 he_srp 6397 70.44 kmr_zh 5042 52.57 pt_te 4191 65.24 fi_fil 3387 57.81 fa_tl 2547 61.07 fil_kmr 2074 56.6
mr_sl 8182 62.8 bs_my 6370 65.88 id_te 5010 64.95 nl_so 4186 51.16 el_si 3353 65.32 fr_tl 2543 62.02 arq_bg 2074 48.1
fr-ca_ka 8179 67.28 bs_hy 6367 72.73 sk_uz 5005 54.18 pt-br_so 4183 49.23 ca_fil 3348 59.76 kk_ta 2543 61.08 bo_cs 2074 41.05
ckb_da 8171 66.99 lv_ta 6361 62.31 eo_gl 4991 63.14 so_tr 4183 49.12 kk_ur 3341 60.06 ko_nn 2540 59.16 bo_fa 2073 44.54
ka_my 8139 60.24 bn_sl 6355 68.08 hu_srp 4988 68.29 hr_srp 4181 69.78 eu_ka 3320 61.51 eo_ur 2540 63.24 gl_te 2068 62.93
et_nb 8124 67.33 sr_te 6311 63.66 el_srp 4987 73.39 hr_sh 4173 70.49 mn_zh 3314 49.6 ru_tl 2539 60.17 eu_uz 2064 52.99
az_hi 8096 60.93 ckb_sl 6304 64.29 de_ne 4983 67.21 az_gl 4169 59.11 sl_sw 3306 62.15 mr_te 2539 60.7 nb_ne 2058 69.48
ml_ro 8049 60.16 ne_sr 6299 69.55 gl_zh 4978 55.63 bs_gl 4161 67.41 nn_vi 3300 68.92 sw_te 2527 63.69 bo_hu 2055 44.54
gl_mr 8007 59.4 srp_zh-cn 6295 63.37 ne_sv 4972 71.52 el_so 4160 48.84 hi_si 3291 69.06 eo_mn 2524 55.01 nn_uk 2045 68.81
te_zh-cn 7956 56.55 mk_mr 6288 58.12 gu_sq 4971 63.2 es_is 4144 63.86 fr-ca_gu 3272 66.58 az_bs 2518 61.94 tl_uk 2041 64.6
kk_sq 7864 56.59 cs_ml 6267 60.7 el_ne 4966 64.92 so_vi $41$ 51.72 ne_sq 3257 62.74 hy_sw 2518 69.15 km_ru 2039 63.8
sv_uz 7852 59.04 eo_hi 6265 64.27 fr-ca_zh 4964 55.22 nl_sh 4132 72.79 is_sr 3237 65.82 bs_ur 2515 70.69 fr-ca_sw 2038 65.18
bn_fi 7841 66.29 kk_lv 6244 59.13 fi_uz 4964 55.96 ro_si 4128 64.58 hi_ne 3236 65.09 arq_ko 2514 52.02 is_mk 2031 63.51
hy_ta 7832 64.75 mk_ta 6231 65.15 nl_srp 4952 71.32 fi_sw 4125 64.04 bs_ckb 3234 66.13 tl_zh-tw 2513 58.91 sq_te 2025 58.79
ckb_nb 7797 66.96 si_tr 6218 64.71 mk_ur 4906 67.02 id_sh 4120 70.69 pt_si 3221 68.91 gu_ur 2512 71.18 et_ne 2024 64.65
hu_te 7792 62.01 eo_hy 6200 66.69 fi_kk 4897 60.4 pl_si 4109 63.66 sh_sv 3179 69.02 kmr_te 2509 59.7 bo_ko 2022 49.64
hr_te 7784 64.98 fil_pl 6193 62.18 de_srp 4877 70.75 de_so 4108 47.21 ar_nn 3159 66.86 bs_mr 2509 62.76 be_bs 2013 65.68
lt_zh 7736 53.52 srp_vi 6184 70.17 be_fi 4871 60.6 id_is 4105 63.45 hy_ne 3159 67.29 arq_fa 2508 53.27 ms_zh 2010 57.47
pt-br_te 7693 64.08 bn_kmr 6178 61.92 lv_zh 4866 54.79 nb_uz 4097 60.78 bn_et 3146 66.46 kn_uk 2505 61.5 el_tl 2005 62.67
bs_mk 7683 69.54 bg_ne 6166 62.05 hy_zh 4856 57.42 is_nl 4096 66.24 az_mr 3144 53.16 arq_tr 2504 51.47 fi_gu 1996 69.12
et_hi 7681 66.08 eu_hy 6164 64.93 be_kmr 4846 59.65 kn_vi 4087 61.01 be_ms 3137 60.12 tl_vi 2501 60.54 bo_id 1985 40.99
az_fi 7640 58.85 fr_ne 6159 66.69 kk_mk 4846 56.7 is_zh-cn 4086 56.53 te_ur 3134 61.74 arq_es 2501 52.93 is_kmr 1984 61.23
my_sw 7636 59.39 de_te 6121 62.09 eo_my 4836 58.46 my_te 4076 60.39 is_sv 3112 66.83 kn_sk 2497 60.15 gu_sl 1981 70.02
hr_ml 7612 64.32 ca_zh 6088 53.32 nb_ur 4833 72.01 si_sr 4073 66.91 arq_zh-tw 3112 52.23 bn_sw 2479 61.97 bn_gu 1980 72.42
mr_nb 7607 64.61 fr-ca_ta 6066 64.43 hi_uz 4787 60.51 ar_kn 4062 59.25 ca_sh 3107 68.67 eo_ms 2474 61.91 hu_km 1967 64.06
sv_sw 7607 70.84 fr_srp 6066 71.8 fr_sh 4770 72.59 lt_ne 4059 63.8 mk_sw 3104 63.23 mr_zh 2464 55.7 es_km 1966 67.72
eo_nb 7569 68.08 ne_zh-cn 6064 62.98 bg_sh 4770 73.21 kn_ru 4050 58.72 da_is 3104 65.84 ar_tl 2457 59.14 bo_pl 1960 43.86
eo_mk 7508 64.2 ne_pt-br 6057 71.48 ko_sh 4770 63.27 ro_so 4049 44.28 arq_ru 3104 52.92 is_ro 2456 63.27 fil_ka 1960 60.52
gu_my 7504 64.04 fil_th 6047 60.1 sh_tr 4770 63.79 da_fil 4045 62.87 is_lt 3098 58.24 be_bn 2451 62.24 km_vi 1956 66.07
en_si 7501 77.36 ca_kk 6045 59.96 sh_zh-tw 4770 62.55 kn_zh-cn 4045 55.27 ar_arq 3092 53.42 lt_sh 2442 63.14 fi_is 1954 60.17
az_ca 7480 55.07 bs_kmr 6036 66.14 it_sh 4770 71.64 hr_is 4042 62.58 hr_so 3088 48.82 kmr_ne 2442 51.84 gl_ne 1953 69.16
ar_ne 7472 60.68 lv_uz 6027 53.48 sh_zh-cn 4770 60.96 fr_kn 4035 62.31 az_eu 3085 52.75 lv_ne 2441 66.83 et_ur 1951 65.06
eo_sl 7465 61.17 ko_srp 6027 64.07 hu_sh 4770 69.83 he_is 4035 62.79 en_sh 3075 79.65 eo_eu 2437 53.92 km_zh-tw 1950 62.36
az_hy 7462 61.64 kmr_sw 6025 51.99 pt-br_sh 4770 72.2 es_kn 4021 63.31 es_nn 3068 66.55 bg_tl 2430 62.53 bs_zh 1943 58.49
bg_fil 7448 62.38 ja_srp 6022 65.21 de_sh 4762 71.85 srp_uk 4019 69.99 eu_hi 3067 63.4 pl_tl 2427 59.46 bn_fil 1930 63.46
ar_fil 7448 58.75 si_zh-cn 6020 62.21 ru_si 4760 62.08 ro_sh 4015 71.22 ms_uz 3063 55.53 nl_nn 2416 68.93 my_si 1928 62.24
fil_it 7447 66 fil_sr 6018 62.44 bg_srp 4760 72.47 kn_pt-br 4011 62.41 my_ne 3062 60.74 et_zh 2415 53.4 arq_sr 1925 57.29
fil_zh-tw 7433 57.01 srp_tr 6009 64.41 gl_mn 4751 56.9 sh_uk 4010 69.58 so_sq 3052 45.33 gu_nb 2413 72.27 kn_te 1921 64.72
ckb_mk 7421 56.04 mn_ur 6000 58.23 bn_mk 4727 66.7 cs_sh 4008 66.87 ne_pt 3048 70.05 de_nn 2412 67.68 fil_uz 1920 49.53
my_ta $73$ 61.97 fi_zh 5993 56.72 sh_vi 4684 69.23 fr-ca_kk 3996 60.73 gu_hy 3027 67.86 az_ms 2406 60.03 ne_sl 1912 68.97
bs_fi 7350 67.02 id_ne 5993 69.43 ckb_ur 4678 54.18 it_kn 3993 62.35 be_gl 3024 62.29 srp_sv 2405 70.83 ja_km 1912 61.08
et_ka 7350 63.07 kmr_uz 5956 51.96 es_sh 4676 72.69 fr-ca_ur 3987 66.84 mr_uz 3011 56.17 kk_zh 2399 56.7 sk_srp 1911 67.44
ka_mn 7335 53.47 mk_ms 5951 66.38 gl_ta 4673 64.32 fr_is 3969 62.71 arq_he 3010 56 kn_ta 2399 61.07 kk_uz 1910 60.78
mn_mr 7334 53.51 mn_ta 5946 57.54 gu_lt 4672 65.2 kk_mn 3965 56.55 be_kk 3001 56.03 sh_sk 2398 68.73 da_sh 1909 69.73
bg_te 7322 61.03 sl_zh 5945 57.5 bg_is 4669 63.47 ckb_sw 3958 51.56 nn_pl 2993 66.77 fi_kn 2397 61.44 bo_pt-br 1907 41.19
eo_pt 7312 65.71 srp_zh-tw 5927 64.8 mn_ms 4668 54.98 gu_mr 3958 64.13 bs_et 2993 65.17 sk_so 2395 44.84 si_sk 1906 65.71
fil_hr 7278 64 fa_srp 5916 70.72 bs_fr-ca 4668 71.63 bn_mr 3945 64.11 ja_kn 2985 59.38 lv_sh 2388 63.01 az_ckb 1888 55.69
Table 14: COMET-QE score and the count of sentences for all 4,765 language pairs in TED2025 (part III).
Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE
fil_sl 1886 59.82 gu_ta 1557 66.39 ja_pa 1304 63.13 hy_srp 1082 65.79 bg_lo 854 64.75 pa_uk 665 62.48 ta_tg 532 44.58
km_ro 1881 67.57 ps_pt 1554 68.04 arq_my 1298 53.16 kn_mn 1082 57.54 eu_lo 854 58.47 si_ur 665 68.2 lo_nl 530 69.82
kn_lt 1878 58.29 th_tl 1553 61.61 mk_nn 1296 70.71 ml_my 1079 59.78 et_lo 854 62.6 ja_sc 664 45.6 af_be 524 59.74
arq_pl 1870 46.11 ca_srp 1552 67.46 id_ug 1296 69.21 tg_uk 1079 37.69 lo_tr 854 62.66 bs_ne 663 67.47 srp_ur 523 66.31
bo_ru 1870 41.9 ms_srp 1549 66.72 af_da 1294 63.71 ht_ko 1076 46.13 lo_th 854 64.58 cs_pa 661 69.44 es_lo 523 65.34
bo_de 1868 41.21 eu_mr 1547 57.91 ms_so 1290 46.86 pa_th 1073 67.43 ar_lo 854 61.37 mg_vi 661 48.29 lo_ru 523 59.54
hi_nn 1865 71.95 pa_zh-tw 1542 65.66 ckb_srp 1283 64.88 gl_nn 1073 65.89 da_lo 854 68.15 fr_mg 661 45.43 eo_kn 522 65.77
ta_uz 1864 60.22 am_de 1542 62.34 de_km 1283 68.27 km_kmr 1072 57.75 it_lo 854 67.64 mg_uk 661 47.49 km_mn 519 53.66
mk_te 1863 60.57 es_pa 1541 70.31 am_ro 1278 57.09 sr_tg 1071 39.21 id_lo 854 69.12 ko_mg 661 47.43 ky_sq 518 57.91
sl_te 1862 63.62 gl_ml 1538 64.06 lv_si 1277 62.72 hi_srp 1067 71.45 ca_lo 854 64.76 es_mg 661 45.23 eo_srp 515 64.9
gu_ms 1856 67.06 he_ps 1537 62.19 pt_so 1276 49.53 am_kmr 1066 59.48 hy_lo 854 63.15 ar_mg 661 44.97 fi_km 514 71
nn_sv 1855 71.61 pa_zh-cn 1529 63.6 bs_fil 1275 61.05 fr-ca_ky 1065 64.75 hu_lo 854 64.05 mg_zh-tw 661 46.56 fi_tg 512 43.93
bs_eu 1847 59.46 az_gu 1526 64.98 et_tl 1274 56.83 lt_tl 1064 54.01 hr_lo 854 66.7 it_mg 661 45.97 bn_so 512 53.43
en_tl 1845 72.59 it_ug 1521 67.45 fr_ht 1274 36.5 fil_my 1063 59.24 lo_nb 854 68.4 id_mg 661 45.08 ka_tg 511 43.56
am_he 1845 59.07 hu_ug 1518 65.35 kmr_ug 1272 52.61 ms_sw 1063 66.49 is_uz 853 53.24 mg_nl 661 51.62 kmr_ps 510 51.24
da_kn 1844 62.59 eo_ne 1512 62.26 he_ug 1271 65.85 hi_pa 1062 73.84 bn_si 845 69.27 he_mg 661 47.03 fi_ky 507 66.41
nn_ro 1843 67.09 af_hu 1510 64.07 ht_tr 1270 43.29 ckb_eo 1059 53.29 fr-ca_nn 843 65.41 ko_sc 661 44.9 az_so 505 46.8
ps_tr 1842 60.51 af_ko 1508 61.4 kmr_srp 1270 64.5 hy_tg 1058 44.51 hi_km 840 66.92 ru_sc 659 40.1 lv_mfe 503 40.07
am_ja 1840 61.1 fil_mk 1507 62.55 ja_ug 1269 64.66 el_pa 1052 66.66 arq_fi 839 53.46 sc_zh-cn 657 42.56 eu_sw 502 55.75
am_tr 1838 62.56 ru_ug 1505 65.89 nl_ug 1269 69.49 da_tl 1050 62.82 lo_zh-tw 839 58.96 mg_sr 655 46.18 ka_ky 501 64.55
bn_zh 1837 59.4 ko_ug 1504 65.48 de_ug 1267 68.1 ta_tl 1049 57.59 gl_so 838 46.95 af_gl 654 63.25 da_ky 501 68.53
am_zh-tw 1835 57.88 be_eo 1504 55.38 ht_zh-cn 1266 41.45 ko_pa 1046 65.11 ky_sr 836 61.85 hy_te 653 62.91 cs_lo 500 64.42
eu_my 1832 59.79 fr_ps 1503 64.41 fa_ht 1263 43.64 arq_sl 1044 43.1 af_ca 831 61.81 et_nn 650 63.89 ps_te 498 65.11
bo_nl 1832 43.99 pa_ru 1501 66.84 gu_sw 1262 68.07 bs_ml 1041 65.4 lo_pt-br 830 68.13 eu_kk 649 56.46 mfe_ru 497 44.48
arq_ro 1831 47.37 nl_ps 1495 66.58 ps_sv 1262 67.79 ne_uz 1038 63.79 am_ca 829 62.28 ro_tg 649 37.67 mr_pa 494 67.28
lt_si 1830 60.89 af_nl 1494 66.21 cs_ug 1260 66.57 hy_tl 1036 65.97 am_hy 829 62.11 bo_mn 649 42.53 tg_uz 493 47.92
km_pl 1830 67.67 bo_lt 1490 42.88 bg_ug 1260 63.15 fi_tl 1036 60.66 kmr_pa 826 60.09 ca_ps 648 61.2 ta_ug 492 64.15
is_sl 1827 62.6 af_ar 1484 63.07 af_sr 1257 67.67 mr_ps 1030 60.12 bo_mk 826 40.06 am_ka 647 60.07 fi_pa 490 69.41
sr_tl 1821 60.05 fr-ca_te 1484 62.77 fil_mn 1256 52.38 ml_sw 1026 61.88 ht_sr 824 39.05 fa_mg 647 49.02 ky_lv 489 63.51
kn_sr 1819 63.54 ps_sr 1483 67.91 mn_te 1254 57.32 si_sq 1022 60.27 sv_ug 824 68.7 bg_mg 647 42.36 fil_si 489 62.48
am_ar 1817 59.12 id_km 1482 70.45 ht_ru 1254 42.48 eu_ml 1022 62.17 ro_ug 823 66.3 mg_tr 647 47.67 et_si 488 62.93
gu_kk 1811 64.14 af_pt-br 1478 68.01 hi_tl 1252 64.59 mn_srp 1022 55.83 ht_nl 823 41.69 de_mg 647 45.47 eu_is 486 53.05
be_zh 1810 50.21 ml_mn 1478 53.69 lv_nn 1247 64.27 fa_ky 1019 64.39 lv_tl 822 57.41 mg_zh-cn 647 42.54 is_ms 486 57.87
it_km 1810 70.53 ps_zh-cn 1476 59.69 nb_srp 1243 73.56 bo_kmr 1019 45.66 hi_ps 822 61.93 mg_pt-br 647 46.99 fr_lo 486 65.89
ca_nn 1804 64.35 am_hu 1476 60.75 be_eu 1242 53.63 ky_uk 1018 63.79 ml_nb 822 65.94 et_te 646 62.95 arq_sh 483 36.33
hy_sh 1804 73.93 af_pl 1472 62.72 de_ps 1241 64.28 et_gu 1015 64.47 eo_tl 821 57.24 kn_mk 646 64.88 eu_sh 483 61.08
lt_nn 1801 61.55 bo_uk 1468 45.21 hi_so 1240 52.92 af_sv 1013 66.86 ne_ur 818 65.05 lv_ug 645 63.04 ckb_ps 482 49.54
fa_km 1801 63.34 bo_el 1468 43.97 eu_ms 1239 56.45 km_my 1008 57.14 km_pt 814 68.93 hy_kn 644 61.5 arq_az 482 45.73
am_vi 1800 63.6 ps_ro 1464 66.71 ht_zh-tw 1236 43.83 arq_gl 1007 37.38 lv_tg 810 40.13 fr-ca_srp 642 68.36 bs_is 478 63.01
gu_ka 1795 66.29 af_sk 1460 64.59 kk_ne 1234 62.26 ps_sq 1006 60.11 srp_zh 806 62.43 sc_tr 640 45.98 lo_vi 478 64.34
am_en 1782 74.17 af_th 1460 65 pa_sr 1227 70.8 lt_ug 1004 61.59 da_tg 805 43.92 be_sw 637 59.2 bs_lo 477 68.19
be_uz 1781 54.17 sv_tl 1459 64.73 fr_pa 1226 69.65 de_ht 1004 36.32 ko_tg 805 45.11 sc_zh-tw 636 46.67 ka_lo 477 59.07
arq_sv 1778 54.62 pa_ro 1458 70.16 hy_nn 1225 67.38 bg_ht 1004 36.77 ht_ja 801 44.38 eu_ur 629 64.06 lo_ta 477 65.35
sq_tl 1771 53.11 th_ug 1454 61.31 bo_hi 1224 46.88 ckb_ne 1000 53.24 lo_zh-cn 801 57.89 nb_ug 628 67.14 de_lo 477 64.62
am_sr 1770 62.3 sr_ug 1452 64.62 he_pa 1218 66.24 arq_hi 999 53.98 kk_te 800 61.77 am_az 628 59.93 hi_lo 477 66.27
az_be 1770 56.56 af_de 1449 64.69 ps_th 1215 60.13 el_km 999 65.89 sl_tg 799 39.38 bo_nb 628 41.73 eu_ne 474 60.86
am_fr 1767 63.61 arq_mk 1445 41.2 ckb_te 1213 58.88 bo_my 998 53.05 am_sl 799 59.13 si_ta 627 64.87 eo_km 474 68.21
bo_ro 1766 36 ar_pa 1444 67.02 pl_ps $12$ 64.09 et_ml 989 58.92 ckb_pa 796 60.95 fr_sc 627 39.22 el_lo 473 62.13
ps_pt-br 1765 67.74 it_ps 1442 66.08 ca_so 1209 45.3 af_sq 987 61.27 nn_ta 795 67.44 ht_uk 626 44.82 ml_ms 468 64.67
ca_te 1756 62.68 arq_th 1441 46.07 ug_uk 1206 63.54 af_ro 985 67.08 my_ps 794 57.27 eo_te 626 57.17 fr-ca_ug 466 66.52
arq_sq 1752 48.41 fr-ca_so 1440 47.35 af_lt 1205 59.86 pt_tl 980 60.33 fil_zh 793 51.9 ht_id 625 41.95 ca_ug 464 65.12
da_so 1751 51.29 es_ps 1439 67.16 ug_zh-cn 1203 62.89 cs_km 980 70.39 kk_sw 791 54.78 gu_uz 625 65.41 ky_sl 462 59.37
fr_km 1739 67.38 af_it 1437 68.37 ht_it 1203 42.06 fr-ca_tg 977 41.57 bo_sk 791 42.44 az_ky 624 59.16 ka_tl 461 62.4
bs_uz 1736 56.61 pa_tr 1437 69.6 ur_uz 1203 56.64 kk_si 976 63.88 am_hi 791 65.52 fa_sc 622 45.43 hr_ht 461 36.67
am_nl $17$ 65.48 af_el 1436 69.8 bo_sq 1202 39.96 de_pa 974 69.18 eu_zh 790 53.29 be_gu 620 61.7 lo_lv 460 65.67
bo_sr 1716 45.9 eo_kk 1433 55.54 bn_ne 1202 68.78 ky_pt-br 969 65.65 ja_lo 789 58.32 sl_ug 619 63.68 si_sl 459 69.34
bn_kk 1714 64.69 af_he 1432 68.21 ky_ro 1199 65.32 sl_tl 963 59.67 gl_lo 786 66.26 hu_sc 619 40.91 gu_ne 458 69.56
am_pl 1713 60.77 bs_kk 1427 65.04 fil_ta 1198 63.48 km_nl 962 70.98 fr-ca_lo 786 63.33 am_ta 618 63.16 pa_sl 458 71.09
mk_ne 1709 61.55 pt_sh 1422 71.43 et_sh 1196 64.42 nl_pa 961 72 lo_sv 786 70.79 bn_tl 617 62.51 ca_ht 457 36.82
mk_sh 1709 70.52 af_hr 1422 70.12 id_pa 1193 71.73 ky_zh-cn 961 57.53 en_tg 785 51.12 is_mr 613 57.27 bo_gl 457 39.88
am_ru 1705 59.86 am_zh-cn 1420 57.84 kmr_si 1187 63.38 ky_sk 960 65.46 gu_zh 784 58.73 mr_sh 612 60.58 ne_tl 455 62.25
am_ko 1697 62.72 hr_km 1416 70.21 pa_pt-br 1186 71.88 en_ht 957 51.36 hy_si 781 67.97 gu_kn 609 65.98 cs_mt 452 42.46
am_id 1695 62.92 mr_ne 1416 60.26 pl_ug 1184 65.24 bs_te 955 66.4 ne_ta 781 69.71 ky_mk 608 58.92 mg_ru 452 43.65
bn_eu 1692 62.28 km_zh-cn 1412 59.62 arq_ca 1174 38.95 ja_ky 952 62.43 af_fi 779 64.58 sv_tg 608 47.03 he_mt 451 48.51
hy_ml 1691 60.19 ml_ta 1408 60.68 km_sv 1172 67.79 lt_tg 950 39.34 nb_tg 771 44.53 gu_te 605 65.17 bs_so 451 47.81
eo_uz 1689 53.17 am_uk 1404 62.72 arq_pt 1168 60.84 is_mn 950 50.15 kmr_so 763 48.06 cs_sc 600 35.49 bo_ta 451 47.38
bo_tr 1685 46.61 am_th 1401 63.01 be_fil 1161 54.31 fil_nn 949 60.41 mr_sw 762 62.05 hy_ug 599 67.04 he_sc 451 42.3
ko_ps 1678 61.93 hi_is 1396 65.18 bn_ml 1161 63.4 cs_tg 948 40.8 am_lt 762 58.98 af_uz 598 53.55 mt_pl 450 44.03
nn_sr 1674 67.74 nb_sw 1388 69.73 hu_ky 1158 63.73 he_ht 945 40.06 bo_eo 760 40.64 bo_fr-ca 598 39.99 mt_zh-cn 450 47.13
fi_srp 1674 67.72 af_cs 1386 65.97 ps_vi 1156 62.06 gl_is 943 62.22 arq_eu 758 38.42 am_pt 597 66.09 sc_uk 450 48.61
sh_sl 1669 66.6 am_pt-br 1384 65.12 af_ja 1155 63.06 bg_ps 940 59.9 id_ps 756 68.34 af_pt 596 66.1 ar_sc 450 41.4
hu_ps 1666 63.79 is_nb 1382 66.48 az_te 1154 61.37 ne_zh 930 58.69 pt_ug 755 67.95 pt-br_sc 593 39.87 ja_mt 449 49.06
bn_bs 1665 71.98 af_bg 1380 68.63 el_ht 1152 41.78 cs_ht 929 32.28 mn_sh 752 55.22 it_sc 592 39.94 sh_uz 447 52.83
am_fa 1665 61.46 af_tr 1380 62.03 be_te 1149 56.58 af_lv 928 64.89 pa_pl 751 70.55 is_kn 592 61.36 fil_sh 447 62.09
ar_ug 1660 59.19 ca_tl 1379 60.17 ht_pl 1147 37.12 be_so 928 43.5 sw_zh 751 58.03 kmr_ky 591 53.29 si_zh 447 56.7
ug_zh-tw 1660 61.29 fa_ug 1374 63.7 tg_vi 1147 36.84 eo_sw 926 57.3 fi_sh 750 65.83 am_mn 591 54.72 mt_ru 447 51.8
uz_zh 1658 52.66 mn_ne 1374 57.76 fa_tg 1147 46.71 it_pa 924 71.78 ka_sh 750 67.04 am_eu 591 59.41 hr_mt 447 46.31
bo_th 1657 45.46 ka_sw 1373 62.95 bg_tg 1147 34.77 is_pt 924 65.64 kn_lv 748 51.62 fi_nn 590 65.98 si_te 447 67.83
fr-ca_sh 1652 72.85 ky_vi 1371 58.37 es_tg 1147 43.71 ckb_ml 921 59.82 srp_ta 747 64.86 kk_srp 590 63.81 mt_zh-tw 446 48.95
arq_da 1648 48.76 fr_ky 1371 64.49 tg_tr 1147 41.64 bo_da 910 40.58 arq_kmr 746 52.68 da_km 589 74.16 ar_mt 445 48.74
bs_sw 1645 68.45 bg_ky 1371 59.94 tg_zh-tw 1147 38.97 az_kn 910 61.81 ms_si 746 67.43 pa_pt 588 73.16 fil_lo 445 59.46
ta_zh 1644 54.63 ko_ky 1371 63.64 ja_tg 1147 44.52 is_ka 907 60.04 fr-ca_is 745 63.38 am_bn 588 62.31 eo_lo 445 62.38
cs_tl 1642 59.34 es_ky 1371 66.14 de_tg 1147 44.69 fr-ca_km 907 68.35 ky_sv 744 66.52 et_tg 587 42.23 az_lo 445 58.94
az_uz 1635 54.26 el_ky 1371 64.43 it_tg 1147 44.48 sk_tl 905 58.5 az_ml 742 58.62 az_km 582 60.88 lo_sl 445 66.15
az_sw 1635 58.46 ky_tr 1371 60.41 tg_zh-cn 1147 38.01 mk_tl 904 61.57 pa_ur 739 71.28 so_ta 581 51.77 fi_ug 444 65.28
af_fa 1634 66.06 ar_ky 1371 60.27 id_tg 1147 46.51 pl_tg 901 40.7 af_ur 734 67.2 srp_sw 580 65.33 am_sq 444 57.07
am_it 1630 65.55 ky_pl 1371 65.44 ru_tg 1147 38.19 ml_mr 901 56.85 ckb_et 731 57.81 hr_pa 578 71.44 pt_tg 443 44.08
en_pa 1626 78.4 ky_zh-tw 1371 60.82 nl_tg 1147 47.71 hy_so 901 51.86 af_hi 727 67.11 am_eo 578 58.71 km_lt 443 60.9
am_cs 1626 61.7 de_ky 1371 66.38 hu_tg 1147 43.51 kn_pt 900 61.02 sk_tg 726 39.71 mg_pl 578 43.98 ka_ps 442 66.14
am_el 1624 60.1 it_ky 1371 66.71 hr_tg 1147 41.14 en_ky 900 75.16 bs_si 724 69.73 ps_sk 574 68.87 hy_ky 442 63.87
am_hr 1623 63.29 id_ky 1371 67.8 he_tg 1145 42.78 el_tg 899 43.19 af_sl 723 66.54 ka_km 574 67.09 bo_mr 441 43.15
am_es 1619 64.18 ky_ru 1371 63.05 el_ps 1144 61.78 ht_vi 895 42.87 arq_lv 722 36.49 gl_sh 574 69.26 kn_zh 439 50.95
am_bg 1615 59.14 hr_ky 1371 64.96 eu_fil 1141 52.22 sq_ug 890 58.68 am_fr-ca 719 63.87 el_mg 573 46.32 kmr_sc 439 46.3
af_vi 1614 66.01 he_ky 1371 64.62 gl_si 1133 64.08 pt-br_tg 890 44.94 ms_tl 716 59.6 es_sc 571 41.15 so_ur 438 53.26
af_zh-cn 1609 59.07 cs_ky 1370 65.34 sl_so 1130 46.09 fil_so 890 45.67 ms_tg 715 46.14 eo_gu 568 66.05 mt_vi 436 49.26
km_uk 1608 66.54 ky_nl 1370 64.32 mk_srp 1130 73.93 mn_si 889 57.8 ckb_eu 711 53.58 mg_th 568 49.16 be_lo 435 52.7
kn_sl 1606 56.72 ug_vi 1369 61.1 fil_ms 1130 57.19 my_ug 888 60.08 is_zh 710 54.23 is_kk 565 56.57 ht_my 434 53.53
es_ug 1605 67.14 be_ckb 1367 61.12 fi_ml 1126 60.7 is_my 887 59.6 ht_hu 704 41.65 bs_ky 565 61.92 am_my 434 61.69
bo_en 1602 33.94 kn_sv 1359 66.37 am_fi 1123 61.42 ca_tg 882 41.28 ca_km 701 68.07 ka_nn 563 63.83 ro_sc 434 31.84
tr_ug 1600 61.99 az_ne 1354 60.68 af_en 1121 75.66 cs_ps 880 65.09 mr_nn 700 64.1 am_sw 563 61.19 sc_th 433 48.06
af_fr 1594 67.14 hr_ug 1349 66.7 arq_sk 1118 40.44 af_kmr 877 62.55 my_tl 699 58.62 pt_sc 563 40.03 bn_srp 431 74.6
hr_ps 1594 64.78 fa_pa 1343 68.64 ur_zh 1117 58.15 ht_ro 874 34.18 my_pa 698 65.91 mk_so 562 46.45 fa_mt 428 46.93
sl_srp 1592 66 km_th 1343 66.37 ht_pt-br 1117 42.98 sq_tg 874 36.38 az_bo 698 44.38 kmr_nn 561 64.53 so_zh 427 46.32
en_ug 1590 73.32 arq_lt 1337 43.99 bs_ms 1116 66.88 ar_tg 874 39.76 arq_nb 697 57.84 ml_ur 560 62.81 mk_tg 427 36.85
af_zh-tw 1586 61.66 nb_te 1330 64.15 bo_fi 1115 46.79 fil_ne 874 63.47 am_da 692 62.12 be_si 560 60.11 hu_mt 426 47.39
fr_ug 1586 65.55 he_km 1327 62.44 kmr_sh 1113 62.99 ka_so 872 49.67 lv_so 687 44.99 eo_nn 558 64.81 nn_so 424 49.97
ka_ml 1586 56.78 km_pt-br 1326 68.81 ca_ky 1109 64.09 hi_tg 871 46.82 hi_ug 682 66.96 ht_pt 556 43.48 bs_sh 422 73.56
af_es 1584 66.06 nb_tl 1324 65.55 bo_sv 1106 44.45 mn_nn 868 54.7 bo_sl 680 43.64 ckb_km 555 52.9 kk_sh 422 65.42
af_ru 1581 65.32 ca_kn 1324 65.84 bo_hy 1103 49.1 az_is 866 54.25 bn_tg 680 48.6 af_ms 554 61.46 nn_uz 420 59.28
ckb_fr-ca 1578 63.68 ar_ht 1319 39.44 kmr_kn 1102 61.25 da_ug 864 64.92 hi_sh 680 71.11 cs_mg 554 38.19 nn_pt 418 68.8
da_srp 1578 72.29 bo_ca 1318 40.25 ky_th 1102 58.63 kn_my 860 61.21 bg_pa 675 65.33 eo_so 553 46.23 el_mfe 418 46.2
ps_uk 1576 61.9 af_id 1314 66.67 fr_tg 1102 41.96 tg_th 860 39.25 hi_ky 675 66.9 gl_ht 551 37.16 fil_tg 418 39.66
lv_te 1574 60.9 pt-br_ug 1314 68.56 gl_tl 1099 60.5 be_srp 860 65.39 arq_hy 674 50.17 da_pa 546 71.22 kn_mr 416 60.66
bo_hr 1569 47.04 nb_si 1314 69.33 hu_pa 1098 67.78 et_sw 855 60.19 gl_srp 673 66.22 fr-ca_kn 546 60.38 fr_mt 415 47.85
fi_si 1569 64.24 es_ht 1314 37.93 sk_ug 1098 64.9 lo_lt 854 60.03 en_sc 673 51.58 eu_nn 542 59.47 et_mg 415 44.79
fil_gl 1568 60.8 mr_si $13$ 64.09 pa_sv 1094 72.32 lo_uz 854 56.6 lt_pa 672 68.09 et_is 542 57.67 id_sc 415 41.82
be_ta 1567 59.29 nb_nn 1312 71.92 be_ne 1091 58.76 fi_lo 854 62.69 is_ta 671 58.8 ms_ne 541 69.67 mr_srp 414 66.22
ka_si 1565 62.76 fr-ca_tl 1312 61.89 pa_vi 1089 69.01 lo_uk 854 65.65 sc_vi 671 45 af_ka 536 66.46 bs_srp 414 72.96
ckb_ka 1559 64.64 km_sr 1306 68.12 bg_km 1086 65.67 fa_lo 854 62.81 pa_ta 668 66.79 he_lo 535 63.53 eu_km 412 66.64
el_ug 1559 65.17 af_uk 1305 67.08 lv_srp 1084 65.01 bn_lo 854 63.09 ms_nn 667 59.84 be_nn 532 64.22 mt_pt-br 411 49.82
Table 15: COMET-QE score and the count of sentences for all 4,765 language pairs in TED2025 (part IV).
Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE
en_mg 411 57.44 id_mt 343 45.08 ga_ru 274 52.87 lb_nl 217 40.42 ug_zh 159 62.63 am_ur 128 66.7 km_sl 94 62.4
km_ky 411 62.79 mn_ug 343 56.64 kn_nb 272 63.34 hu_lb 217 38.76 tl_ug 159 63 ceb_uk 128 48.32 ckb_is 93 61.37
eu_ky 411 59.86 tg_tl 342 38.36 ga_ro 271 50.93 hr_lb 217 39.3 hy_pa 158 60.44 lv_ps 128 61.37 ckb_fil 93 53.04
eo_ky 411 60.48 en_mt 340 64.14 km_sk 271 70.12 he_lb 217 38.13 el_tt 157 55.9 ig_pt 127 38.76 fo_my 93 50.27
am_ky 411 59.28 et_mt 340 42.85 bn_ps 271 64.38 lb_pt-br 217 41.76 cs_tt 156 57.62 mg_pt 125 49.39 fr_ig 92 36
ky_ta 411 68.26 fr-ca_mt 340 45.25 ar_ga 270 51.87 lb_my 217 49.83 fr_tt 156 55.26 am_ms 125 60.65 fo_fr 92 51.41
am_km 411 60.48 da_mt 340 46.18 ga_my 269 53.86 te_uz 217 59.9 fa_tt 156 59.29 af_bo 122 36.83 fo_sr 92 48.98
km_ta 411 69.29 ca_mt 340 46.21 kmr_tg 269 44.45 ky_ur 213 62.34 bg_tt 156 51.89 mg_ur 121 50.09 ky_zh 92 58.8
eo_tg 410 39.13 gl_ry 340 46.18 am_te 268 65.42 bn_ky 213 63.36 es_tt 156 57.47 ms_tt 121 63.59 cs_ig 91 33.64
el_sc 410 43.82 ry_vi 340 43.83 af_my 266 64.68 be_ky 213 55.38 en_tt 156 63.54 lo_ro 121 69.79 el_ig 91 40.48
am_sv 409 65.33 cs_ry 340 45.68 fil_ug 266 58.62 ky_te 213 57.81 th_tt 156 55.24 ckb_kn 120 58.58 fr-ca_tt 91 49.6
lo_pl 409 64.39 lv_ry 340 43.11 kk_tg 264 44.78 af_az 212 61.65 pl_tt 156 57.92 ckb_ig 120 44.92 nn_tl 91 56
lo_sr 409 72.54 lt_ry 340 42.06 gl_km 264 64.79 ga_sk 210 52.77 tt_zh-tw 156 55.94 fi_ig 120 40.05 bo_ml 91 44.72
lo_sq 409 65.53 fr_ry 340 42.75 arq_gu 260 69.16 tr_tt 209 55.3 sr_tt 156 57.84 ig_th 120 45.33 mt_zh 91 45.93
az_tg 408 46.01 fi_ry 340 45.17 ja_tt 260 55.81 tl_ur 209 63.2 sq_tt 156 49.28 hy_ig 120 45.74 gu_ht 90 56.37
it_mt 407 50.87 ry_uk 340 46.24 tt_zh-cn 260 50.97 af_fr-ca 208 67.69 sl_tt 156 56.32 hr_ig 120 38.77 fo_pt-br 90 51.59
ne_te 407 65.71 fa_ry 340 42.65 is_ne 259 62.19 bn_bo 208 48.14 sk_tt 156 57.21 bo_et 120 42.86 is_tg 89 40.67
km_sq 404 61.16 bg_ry 340 41.63 am_gl 259 63.35 ast_nl 208 60.6 it_tt 156 56.32 sk_tlh 120 29.16 bo_ky 89 48.07
km_mr 404 54.45 ko_ry 340 42 ms_te 258 63.71 ky_pt 207 61.39 ru_tt 156 55.68 he_mfe 119 42.25 mfe_uk 87 46.78
nn_sl 403 66.81 es_ry 340 45.8 ne_ps 258 60.89 en_ig 206 45.45 ca_tt 156 56.2 ig_mr 119 44.55 mfe_sl 87 38.79
km_lv 403 65.04 en_ry 340 52.71 tt_uk 256 50.56 ar_ig 206 36.14 hy_tt 156 56.95 ca_ig 118 38.67 fa_fo 86 50.82
arq_ps 403 54.75 el_ry 340 45.08 mn_pa 252 60.37 srp_uz 206 56.87 nl_tt 156 61.72 cs_dz 118 38 fo_zh-tw 86 45.64
ml_so 400 50.8 ry_tr 340 44.06 mn_tl 251 52.44 af_mr 205 62.61 hu_tt 156 57.34 dz_ko 118 46.66 ckb_ky 85 55.71
gu_si 399 68.7 az_ry 340 43.89 fil_gu 249 65.46 mt_uk 204 51.74 hr_tt 156 58.33 dz_en 118 34.21 ht_sk 84 34.08
mg_ro 397 44.07 ry_th 340 43.91 gu_tl 247 64.21 ga_pt 203 55.8 nb_tt 156 60.72 dz_ru 118 42.23 af_fil 84 56.65
eo_is 396 59.56 ar_ry 340 42.87 mt_nl 247 50.1 ceb_vi 203 47.19 he_tt 156 57.51 dz_ne 118 49.27 te_zh 84 55.34
gu_is 394 64.31 pl_ry 340 44.55 af_mn 246 60.52 ceb_cs 203 45.08 bn_km 156 62.32 ar_dz 117 38.43 am_srp 83 64.36
es_mt 394 46.83 ry_zh-tw 340 37.97 ht_sv 245 40.18 ceb_lv 203 44.33 bo_ckb 155 45.43 dz_sv 117 45.4 be_tl 83 62.58
eo_ug 392 58.28 ja_ry 340 40.97 arq_bn 245 60.64 ceb_fr 203 45.79 bo_ka 155 42.72 dz_sr 117 46.52 aeb_es 82 74.04
ky_lt 392 61.83 ry_sr 340 50.85 si_sw 244 65.21 ceb_fa 203 49.85 sc_ur 155 53.7 da_dz 117 38.45 aeb_sr 82 75.91
fr-ca_mg 390 46.06 de_ry 340 44.73 fa_ga 243 55.32 bg_ceb 203 43.82 fr_mfe 154 43.14 dz_uk 116 45.61 ht_sl 81 42.05
kmr_tl 386 54.59 ry_sl 340 44.93 bg_ga 243 54.46 ceb_es 203 45.42 kk_ps 154 59.21 dz_ja 116 47.77 mfe_nl 80 46.56
sw_uz 386 56.5 ry_sk 340 49.52 ga_ja 243 52.08 ceb_en 203 60.16 ps_zh 154 56.72 dz_he 116 42.82 bg_ig 80 37.05
nb_so 385 51.03 da_ry 340 46.1 de_ga 243 54.13 ceb_th 203 48.41 ca_pa 153 73.6 aeb_gu 115 70.39 ceb_tr 80 44.6
bo_lv 384 43.53 it_ry 340 38.59 ga_nl 243 58.71 ar_ceb 203 43.8 he_ig 152 37.36 aeb_ckb 115 60.5 af_ne 80 66.31
ne_tg 384 50.62 ry_zh-cn 340 35.72 ga_hu 243 54.85 ceb_fr-ca 203 45.62 da_ga 152 57.59 aeb_ko 115 64.51 mk_tt 80 53.69
bn_sh 384 67.99 id_ry 340 45.95 ga_hr 243 58.83 ceb_pl 203 44.98 da_ps 151 70.4 aeb_en 115 79.57 lv_tt 80 60.86
my_sh 384 61.17 ru_ry 340 40.11 en_mfe 240 53.77 ceb_zh-tw 203 45.35 ga_uk 150 57.4 aeb_ru 115 68.07 lt_tt 80 60.45
arq_zh 382 50.92 ro_ry 340 40.86 bs_gu 240 71.66 ceb_ja 203 49.31 ig_ja 150 45.72 ig_pl 114 37.2 tt_uz 80 62.09
mt_tr 381 47.02 nl_ry 340 47.36 am_si 240 64.09 ceb_it 203 48.23 az_si 150 63.08 id_ig 114 37.23 bn_tt 80 65.96
de_sc 381 40.71 hu_ry 340 46.23 el_ga 239 54.9 ceb_zh-cn 203 41.58 lt_mfe 149 37.1 pt-br_tt 114 56.5 et_tt 80 60.03
arq_fil 380 39.67 hr_ry 340 47.22 my_tg 239 44.19 ceb_id 203 43.73 fa_mfe 149 44.53 bn_pa 114 73.42 az_tt 80 63.77
arq_eo 380 39.41 he_ry 340 43.4 et_ug 238 61.05 ceb_ru 203 44.83 bg_mfe 149 38.78 pa_ps 114 65.05 tg_tt 80 46.58
fil_ml 380 62.09 pt-br_ry 340 48.11 arq_bs 237 38.83 ceb_hu 203 46.77 mfe_tr 149 44.25 ig_nl 113 45.37 ne_tt 80 63.62
bs_ps 379 66.01 af_nb 339 66.76 arq_nn 237 37.52 ceb_hr 203 45.24 mfe_th 149 50.1 ro_tt 113 53.92 mfe_ro 79 35.24
az_srp 378 61.13 kk_so 336 48.94 arq_ne 237 48.96 ceb_hi 203 51.02 ar_mfe 149 40.85 dz_ro 113 38.39 hr_mfe 79 37
hr_mg 377 43.62 mk_si 334 68.9 am_nb 235 62.83 ceb_he 203 47.07 da_mfe 149 37.5 ckb_mt 112 50.58 gu_tg 79 55.23
ps_si 377 69.01 bo_ms 334 41.23 tt_vi 235 50.8 nl_sc 202 41.68 ps_ur 149 65.01 mt_my 112 57.45 gl_mg 79 46.68
gl_inh 377 27.65 ky_nb 333 63.99 bo_kk 235 48.64 ig_ro 201 30.88 ceb_fil 149 45.63 da_tt 112 62.21 fi_mg 79 45
inh_lt 377 28.05 ht_th 332 45.94 af_et 233 63.26 ast_ro 200 52.11 ceb_ko 149 47.74 dz_vi 112 42.32 bs_mg 79 45.46
inh_lo 377 32.15 ml_te 331 61.03 sv_tt 232 59.3 ceb_el 199 47.71 ceb_sr 149 45.58 dz_nl 112 44.42 mg_tl 79 47.87
lo_ne 377 61.77 bo_uz 331 45.42 ast_gl 232 56.35 am_et 199 60.38 ceb_sq 149 39.11 aeb_ar 112 66.52 mg_sl 79 47.35
inh_uz 377 33.12 bo_fil 331 41.75 ast_vi 232 55.05 am_kk 198 59.38 ceb_de 149 44.79 dz_tr 111 46.89 mg_nb 79 49.89
fi_inh 377 30.2 ne_nn 328 70.52 ast_cs 232 55.24 ml_tg 197 44.26 ms_ps 149 63.6 aeb_zh-tw 111 64.53 lb_mk 79 44.67
inh_uk 377 29.9 az_tl 327 56.67 ast_lt 232 53.25 so_tg 197 40.17 mfe_zh-tw 148 47.82 bn_ug 110 68.13 gl_lb 79 36.54
fa_inh 377 31.56 az_ug 326 60.88 ast_fr 232 57.55 sc_sr 197 48.52 mfe_nb 148 38.32 dz_it 110 43.01 lb_lv 79 40.75
bn_inh 377 38.56 tg_zh 323 34.63 ast_uk 232 58.02 cs_mfe 196 36.38 lt_ps 148 57.12 dz_es 109 41.8 lb_lt 79 39.68
bg_inh 377 25.27 bo_zh 322 37.56 ast_fa 232 55.09 ig_pt-br 195 37.5 mfe_zh-cn 147 43.99 it_mfe 108 40.37 fi_lb 79 40.46
be_inh 377 28.44 eu_te 318 59.75 ast_bn 232 60.32 gu_pa 195 70 kk_kn 147 60.39 ckb_so 108 49.4 kk_lb 79 44.8
fil_inh 377 28.73 hr_sc 318 41.12 ast_bg 232 60.4 af_eo 194 64.08 be_mfe 146 40.06 dz_th 107 47.3 lb_ta 79 48.63
eu_inh 377 30.93 am_mk 317 62.2 ast_es 232 57.94 ckb_ug 194 52.17 eu_mfe 146 40.6 hu_ig 106 40.65 fr-ca_lb 79 36.14
et_inh 377 29.38 ps_sl 317 69.41 ast_en 232 70.13 mt_ro 194 44.74 ig_it 146 39.85 be_mg 106 42.42 lb_sq 79 40.13
eo_inh 377 29.37 lv_pa 313 67.57 ast_tr 232 53.75 ig_tr 194 42.21 ig_zh-cn 146 37.64 mg_sw 106 44.22 lb_sl 79 40.17
inh_tr 377 32.09 arq_mr 312 55.92 ast_th 232 53.79 ast_pt-br 194 60.14 hu_mfe 145 39.78 az_mg 105 46.8 is_lb 79 38.05
az_inh 377 32.66 pa_sq 311 63.1 ar_ast 232 48.66 sw_tl 194 54.14 bo_pt 145 41.4 gl_mt 105 49.66 ca_lb 79 37.12
inh_th 377 30.35 kk_ky 309 60.42 ast_pl 232 53.55 km_zh 191 53.4 gl_mfe 144 40.11 arq_ckb 104 60.96 lb_ne 79 49.69
ar_inh 377 29.46 ky_mr 309 58.43 ast_zh-tw 232 50.14 eo_ml 191 62.03 ig_sq 143 35.87 mg_my 103 54.91 bn_mg 78 49.17
fr-ca_inh 377 27.67 my_nn 308 61.43 ast_sl 232 52.18 ca_mg 188 45.61 arq_tl 143 39.45 nb_pa 103 76.77 aeb_fr 78 71.85
inh_sv 377 28.27 arq_is 306 60.4 ast_it 232 57.59 da_mg 185 48.8 arq_ta 143 49.54 mfe_vi 102 51.51 fi_ht 77 38.82
inh_ja 377 30.44 bo_kn 306 51.52 ast_zh-cn 232 47.15 hy_mg 185 51.58 arq_ms 143 40.26 am_uz 102 59.89 mg_uz 77 43.85
inh_sl 377 26.05 bo_is 306 43.21 ast_id 232 55.63 hi_mg 185 53.86 so_sw 143 47.54 ast_ja 102 51.86 aeb_pt-br 77 74.14
da_inh 377 28.13 en_ga 304 68.22 ast_ru 232 52.35 bo_sw 185 44.89 km_nb 141 61.07 pa_te 102 68.76 ar_tt 76 45.36
inh_it 377 26.73 ga_zh-tw 303 51.99 ast_hu 232 56.45 hy_km 184 61.64 lb_pt 141 41.58 aeb_hr 102 72.22 kmr_tt 76 49.16
inh_zh-cn 377 26.33 ga_zh-cn 303 49.25 ast_hr 232 56.67 af_bn 183 65.58 mt_pt 141 51.26 arq_ur 102 65.66 ml_srp 76 66.15
id_inh 377 28.85 ga_he 303 54.93 ast_he 232 55.82 fa_ig 181 41.02 mfe_sv 140 43.63 gu_mfe 101 52.52 srp_tl 76 59.19
ca_inh 377 26.94 hu_mg 303 43.26 arq_mn 232 41.4 mr_tg 181 41.4 ja_mfe 138 45.53 mfe_ur 101 53.06 arq_ug 76 55.14
hy_inh 377 32.01 ga_id 302 56.95 el_mt 231 49.28 si_uz 180 64.18 bs_tl 137 63.54 kmr_mfe 101 45.36 be_ug 76 64.97
hu_inh 377 30.1 sc_sv 302 46.72 km_te 231 62.42 ig_ru 179 38.62 az_ps 137 59.48 mfe_my 101 52.07 tg_ug 76 50.13
hr_inh 377 27.59 fil_ur 301 64.54 ga_sr 230 58.06 ceb_nl 179 49.25 ps_so 137 50.25 fo_uk 101 53.39 arq_be 76 43.21
inh_ne 377 38.72 kn_sq 301 58.23 hi_tt 230 55.85 mg_sq 178 42.55 ps_ta 137 59.72 en_fo 101 57.47 arq_tg 76 37.82
inh_nb 377 26.69 ka_pa 300 68.15 lv_mt 229 43.22 ast_mk 178 57.41 km_si 135 60.51 fo_th 101 52.92 be_tg 76 36.26
inh_pt-br 377 27.07 bs_pa 300 73.7 fil_te 229 54.81 ast_uz 178 49.86 my_tt 135 54.44 fo_zh-cn 101 41.83 be_is 75 60.8
bo_gu 375 60.08 ml_sh 300 59.66 ast_el 228 60.2 ast_ko 178 50.94 mn_ps 135 51.3 fo_ru 101 48.99 fo_tr 75 47.99
ckb_ht 373 45.09 km_ur 300 62.66 mr_mt 228 47.63 ast_ka 178 58 ka_ug 135 61.13 fo_hu 101 48.89 aeb_ro 75 74.22
fr-ca_si 371 69.88 ga_vi 299 52.71 af_ta 227 66.64 ast_sv 178 59.09 mr_so 134 46.78 aeb_zh-cn 101 62.04 kk_ml 75 67.04
gu_so 370 54.91 ga_ko 299 56.29 af_hy 227 69.54 ast_sr 178 60.67 fr-ca_ps 132 63.17 fo_ko 100 50.24 ht_ur 74 50.61
kk_tl 368 56.95 tl_zh 297 51.92 ca_si 227 67.09 ast_sq 178 51.53 id_mfe 131 39.92 ar_fo 100 47.58 mr_tt 74 48.69
da_si 368 63.06 fr_ga 297 52.17 mr_sc 226 46.11 ast_de 178 55.96 ps_sw 131 61.15 fo_ja 100 52.37 pa_uz 74 53.96
ht_sq 368 37.2 mn_so 297 43.49 lt_sc 223 37.76 ht_sw 177 39.68 de_ig 131 37.02 de_fo 100 46.71 hi_ig 73 46.26
ckb_tl 367 56.26 ht_kmr 296 51.27 de_tt 222 58 ky_my 176 63.25 ms_ug 131 70.85 fo_sk 100 43.47 pa_zh 73 58.59
ht_ms 367 38.16 ga_pl 296 53.7 hi_sc 222 50.88 eo_ht 176 38.47 ml_tlh 131 38.96 fo_it 100 51.21 en_lo 73 75.68
ka_srp 365 69.16 sh_sw 296 68.36 mr_tl 221 57.56 ml_uz 175 56.95 tlh_vi 131 34.48 fo_he 100 49.67 ug_uz 73 55.71
fi_ps 365 63.2 ga_pt-br 293 56.66 fil_mr 221 53.41 lv_mg 174 43.41 cs_tlh 131 29.24 ug_ur 100 62.87 kn_ms 71 58.99
fi_so 364 45.05 ga_tr 292 51.91 arq_ka 220 63.26 ast_mr 172 45.32 fr_tlh 131 30.11 aeb_tr 100 64.82 dz_pt-br 71 42.79
my_so 364 52.2 ne_si 290 69.61 lb_vi 217 41.73 ko_tt 171 60.91 fi_tlh 131 37.23 kn_uz 100 58.66 si_tl 71 59.77
ja_mg 363 47.34 ky_mn 290 55.05 cs_lb 217 33.56 ceb_ro 171 40.35 tlh_uk 131 37.85 af_ckb 99 66.75 fo_kmr 70 52.37
bs_nn 363 65.01 mk_pa 289 67.71 fr_lb 217 33.96 kk_ug 170 55.6 fa_tlh 131 37.58 es_fo 99 47.47 gu_srp 69 71.68
az_nn 363 59.39 am_sk 289 64.69 lb_uk 217 44.94 ig_vi 169 44.19 bg_tlh 131 31.03 km_ms 99 62.75 bo_srp 69 46.86
bo_ne 363 51.01 am_lv 288 57.38 fa_lb 217 40.64 arq_so 169 44.12 ko_tlh 131 40.21 cs_ga 99 53.15 ckb_fo 69 47.43
inh_zh-tw 362 27.68 bn_nn 288 66.52 bg_lb 217 37.7 ig_ko 168 46.07 es_tlh 131 31.23 ga_lt 99 52.41 kn_si 68 67.76
ht_lt 361 37.39 ht_mk 287 41.97 ko_lb 217 40.7 ml_ne 166 62.88 eo_tlh 131 31.94 fi_ga 99 57.59 en_tk 68 50.49
ka_kn 360 62.93 kn_ml 286 60.01 es_lb 217 37.09 ca_sc 166 39.65 tlh_tr 131 38.59 et_ga 99 50.2 he_tk 68 41.19
kn_ne 360 65.35 is_ml 286 56.01 en_lb 217 41.64 id_tt 165 61.84 th_tlh 131 38.65 ga_th 99 56.1 km_mk 68 57.04
gl_tg 359 41.39 is_ur 286 60.78 el_lb 217 39.93 ceb_pt-br 165 46.38 ar_tlh 131 33.49 ga_sq 99 49.84 af_eu 68 57.01
ko_mt 358 48.12 bs_kn 286 67.94 lb_tr 217 43.95 hi_lb 165 45.4 pt_tlh 131 30.68 ga_hy 99 59.56 tg_ur 68 48.18
fil_kk 355 57.93 bn_is 286 58.97 lb_th 217 43.87 be_km 162 56.21 pl_tlh 131 30.49 ga_ne 99 56.47 eu_tg 68 42.6
ko_lo 353 58.26 is_te 286 58.14 ar_lb 217 37.84 gu_nn 162 67.16 tlh_zh-tw 131 36.41 ga_mr 99 52.18 lo_nn 68 59.81
gu_ml 352 67.45 is_sw 286 64.28 lb_pl 217 38.39 nn_ur 162 69.14 ja_tlh 131 39.98 be_ml 99 57.71 ko_tk 67 44.57
arq_et 352 40.9 ga_it 285 60.43 lb_zh-tw 217 40.92 fil_kn 162 64.31 sr_tlh 131 36.58 lt_mg 99 37.87 et_so 67 44.84
tl_uz 350 51.27 nb_ps 283 67.26 lb_sv 217 41.08 fil_sw 162 55.7 sq_tlh 131 28.26 mfe_sr 98 40.46 so_tl 67 48.15
ka_te 348 63.76 mg_sk 282 41.18 ja_lb 217 39.17 kn_nn 162 62.97 de_tlh 131 30.37 ka_mfe 97 43.17 gl_pa 67 74.88
mk_ug 347 62.04 et_srp 280 66.42 lb_sr 217 44.6 nn_sw 162 64.62 it_tlh 131 31.21 mfe_mk 96 45.72 eu_srp 67 57.8
ml_zh 346 52.75 mn_tg 279 37.96 de_lb 217 36.51 es_ig 161 35.88 tlh_zh-cn 131 33.03 ky_uz 96 64.83 de_mfe 66 37.17
arq_fr-ca 345 39.14 fil_tl 278 67.26 lb_sk 217 38.49 am_zh 161 51.27 id_tlh 131 31.76 ky_ms 96 69.19 af_kk 66 62.77
lt_mt 343 41.06 mg_sv 278 48.68 da_lb 217 35.25 bo_ur 161 45.7 ru_tlh 131 32.78 af_bs 96 68.18 ar_tk 66 39.23
bg_mt 343 44.82 gl_ky 277 65.72 it_lb 217 36.58 mr_ug 161 55.97 nl_tlh 131 32.88 el_fo 96 50.09 pt-br_tk 66 41.5
mt_sv 343 50.4 es_ga 277 51.59 lb_zh-cn 217 38.22 lb_ro 160 36.71 hu_tlh 131 35.67 fo_ro 96 43.85 fo_pl 66 48.09
mt_sr 343 46.31 af_mk 276 64.52 id_lb 217 36.52 ht_ta 160 48.56 hr_tlh 131 34.67 ig_zh-tw 95 40.85 eu_tl 66 54.91
de_mt 343 47.63 is_nn 275 65.05 lb_ru 217 40.38 ckb_nn 159 67.29 nb_tlh 131 32.3 am_ne 95 66.39 eu_so 66 48.28
mt_sk 343 47.24 eu_gu 274 63.07 hy_lb 217 44.88 gl_ug 159 69.93 pt-br_tlh 131 31.2 am_mr 95 64.7 pa_si 66 77.97
Table 16: COMET-QE score and the count of sentences for all 4,765 language pairs in TED2025 (part V).
Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE Lang Pair #Sents COMET-QE
bo_ga 66 39.42 as_uz 57 56.78 ast_eo 54 56.27 bi_fr 49 38.21 gl_la 33 46.56 ro_szl 23 38.62 ne_sw 4 77.92
ga_sv 66 62.57 as_fi 57 68.52 eo_oc 54 37.63 bi_fi 49 39.17 ga_lv 33 54.87 kmr_mg 22 47.65 gu_hup 3 39.39
ga_sl 66 53.98 as_uk 57 64.42 en_oc 54 45.12 bi_uk 49 43.11 ga_la 33 46.36 oc_ro 22 30.58 gu_mt 3 56.14
szl_vi 66 37.79 as_bg 57 63.52 oc_tr 54 42.44 bi_fa 49 45.28 ga_kk 33 50.93 lv_tlh 22 34.58 ckb_hup 3 38.44
cs_szl 66 34.42 as_ko 57 63.2 ast_az 54 53.9 bg_bi 49 33.37 eo_ga 33 53.9 ka_tlh 22 39.07 hup_vi 3 36.74
lt_szl 66 37.71 as_es 57 67.51 az_oc 54 41.61 bi_ko 49 48.18 ga_tl 33 48.45 ig_ta 22 38.14 cs_hup 3 29.79
lo_szl 66 47.98 as_en 57 77.97 oc_th 54 42.46 bi_kk 49 48.99 ga_ta 33 58.61 ha_ro 22 48.38 hup_lv 3 33.43
fr_szl 66 36.82 as_el 57 66.17 ar_oc 54 36.47 bi_et 49 38.02 fr-ca_ga 33 55.72 mfe_sk 21 31.81 hup_lt 3 30.8
szl_uz 66 42.59 as_tr 57 55.52 am_ast 54 51.36 bi_es 49 38.23 ca_ga 33 56.03 af_so 21 45.13 fr_hup 3 31.36
fi_szl 66 40.53 as_th 57 57.17 am_oc 54 43.03 bi_en 49 51.24 ga_nn 33 60.46 mfe_mn 20 46.75 hup_uk 3 31.49
szl_uk 66 40.47 ar_as 57 57.84 ast_ta 54 55.77 bi_el 49 41.61 ga_nb 33 60.59 ga_zh 20 42.99 fa_hup 3 32.02
fa_szl 66 41.86 as_ta 57 64.43 oc_ta 54 45.54 bi_ka 49 46.65 ga_hi 33 59.97 af_sw 20 69.21 bg_hup 3 33.05
bs_szl 66 39.55 as_pl 57 69.9 ast_fr-ca 54 59.54 bi_tr 49 42.66 la_vi 33 53.59 af_is 18 60.57 hup_ko 3 30.66
bn_szl 66 50.47 as_zh-tw 57 61.33 fr-ca_oc 54 36.55 az_bi 49 41.84 cs_la 33 44.39 am_fil 18 55.87 eu_hup 3 32.84
bg_szl 66 39.14 as_sv 57 70.23 af_ast 54 53.37 bi_th 49 49.53 la_lv 33 47.96 am_tl 18 55.66 eu_mt 3 54.81
fil_szl 66 41.31 as_ja 57 65.09 af_oc 54 35.75 ar_bi 49 37.41 la_lt 33 48 af_tg 18 47.77 hup_kk 3 37.83
eu_szl 66 42.46 as_sr 57 72.55 oc_pl 54 37.4 am_bi 49 46.75 fr_la 33 43.81 kk_pa 18 68.09 kk_mt 3 43.48
et_szl 66 41.4 as_sq 57 58.7 oc_zh-tw 54 40.32 bi_fr-ca 49 38.44 fi_la 33 51.61 pt_tk 17 38.63 es_hup 3 30.86
es_szl 66 38.67 as_de 57 66.92 ast_da 54 59.53 bi_pl 49 33.5 la_uk 33 54.17 pt_tt 17 53.84 en_hup 3 29.72
eo_szl 66 39.76 as_sl 57 68.44 ast_oc 54 33.37 bi_zh-tw 49 45.55 fa_la 33 51.22 bi_id 17 36.83 el_hup 3 33.82
ka_szl 66 41.77 as_sk 57 71.59 ast_ca 54 55.96 bi_sv 49 37.58 bg_la 33 47.54 af_pa 17 69.31 hup_ka 3 31.49
szl_tr 66 42.14 as_da 57 71.77 ast_hy 54 61.22 bi_ja 49 45.92 ko_la 33 46.07 gu_ps 17 79.79 ka_mt 3 50.52
az_szl 66 45.07 as_it 57 71.19 ast_nb 54 57.88 bi_sr 49 34.43 kk_la 33 51.07 ps_uz 17 55.63 hup_tr 3 36.8
szl_th 66 41.9 as_zh-cn 57 58.96 ast_hi 54 58.87 bi_sq 49 29.52 et_la 33 48.72 kk_lo 17 54.79 az_hup 3 38.06
ar_szl 66 38.4 as_id 57 69.72 ja_oc 54 39.1 bi_de 49 35.74 es_la 33 47 hi_ht 16 45.61 az_mt 3 55.83
szl_ta 66 45.4 as_ru 57 67.82 oc_sl 54 41.92 bi_sk 49 31.76 eo_la 33 48.21 oc_pt-br 16 35.98 hup_th 3 32.44
szl_zh-tw 66 39.07 as_ca 57 66.63 da_oc 54 34.6 bi_it 49 36.19 en_la 33 55.51 inh_pt 16 25.43 mt_th 3 70.32
de_szl 66 38.3 as_nl 57 71.8 it_oc 54 34.72 bi_zh-cn 49 37.64 la_tr 33 50.01 kk_szl 16 51.57 ar_hup 3 32.26
sl_szl 66 37.61 as_hu 57 63.94 oc_zh-cn 54 36.78 bi_kmr 49 48.81 la_tl 33 49.54 en_szl 16 48.04 hup_pl 3 28.72
da_szl 66 37.86 as_hr 57 71.5 id_oc 54 35.65 bi_ca 49 35.08 la_th 33 53.47 el_szl 16 46.91 hup_zh-tw 3 39.56
it_szl 66 37.53 as_nb 57 71.47 oc_ru 54 39.8 bi_hy 49 51.47 la_ta 33 57.38 aeb_nl 15 73.01 hup_sv 3 33.16
szl_zh-cn 66 36.88 as_he 57 63.5 ca_oc 54 36.06 bi_nl 49 40.14 fr-ca_la 33 45.14 ig_uk 14 42.13 hup_ja 3 30.64
id_szl 66 38.52 as_pt-br 57 70.51 hy_oc 54 43.45 bi_hu 49 40.15 la_pl 33 49.28 el_la 14 46.06 hup_sr 3 31.32
ru_szl 66 37.62 as_my 57 57.65 hu_oc 54 37.66 bi_hr 49 35.34 la_zh-tw 33 48.08 ht_uz 14 35.97 hup_sq 3 35.81
ca_szl 66 38.23 af_am 57 60.76 hr_oc 54 37.86 bi_nb 49 35.78 ja_la 33 47.59 be_ps 14 68.61 mt_sq 3 62.59
hy_szl 66 46.79 te_tl 57 56.02 nb_oc 54 34.22 bi_hi 49 46.17 la_sr 33 54.11 ht_kk 14 46.7 de_hup 3 28.75
nn_szl 66 36.59 mg_zh 56 40.35 hi_oc 54 44.58 bi_he 49 40.82 la_sq 33 50.85 ht_zh 14 35.65 hup_sl 3 28.28
nl_szl 66 40.73 be_szl 56 37.63 he_oc 54 37.05 bi_pt-br 49 37.58 de_la 33 45.85 ht_si 14 42.03 mt_sl 3 71.04
hu_szl 66 42.47 ast_lv 55 53.6 fo_lt 53 40.98 bi_my 48 54.55 da_la 33 46.43 sv_tlh 14 32.44 hup_sk 3 32.23
hr_szl 66 39.6 as_hi 54 58.57 inh_nl 53 29 inh_lv 48 28.13 it_la 33 46.1 af_te 13 65.93 hup_it 3 29.7
nb_szl 66 38.78 ceb_gl 54 41.93 as_fa 52 58.57 lo_mfe 48 49.16 la_zh-cn 33 49 en_inh 13 33.91 hup_zh-cn 3 29.81
hi_szl 66 47.68 gl_oc 54 34.8 as_mr 52 54.57 mfe_uz 48 43.31 id_la 33 46.14 pa_sk 12 74.26 hup_kmr 3 36.9
he_szl 66 39.77 oc_vi 54 38.14 fo_pt 52 49.06 fi_mfe 48 40.34 la_ru 33 50.28 ht_lv 11 33.61 kmr_mt 3 58.67
tk_vi 65 42.63 cs_oc 54 31.64 te_ug 52 62.42 bn_mfe 48 49.83 la_ro 33 44.44 az_pa 11 63.17 hup_id 3 36.06
ru_tk 65 38.16 lv_oc 54 34.48 lo_pt 52 63.98 fil_mfe 48 39.56 ca_la 33 47.97 pa_so 11 55.94 hup_ru 3 32.72
mg_mk 65 36.82 ceb_lt 54 41.54 ha_vi 51 50.3 et_mfe 48 37.96 hy_la 33 52.77 arq_te 11 44.83 hup_ro 3 29.96
eo_mg 65 39.07 lt_oc 54 33.55 cs_ha 51 41.48 eo_mfe 48 38.64 la_nn 33 50.38 am_arq 11 45.79 hup_nl 3 29.42
mg_ta 65 46.38 fr_oc 54 36.58 fr_ha 51 45.08 az_mfe 48 41.38 la_nl 33 51.85 so_te 11 52.72 hu_hup 3 25.78
ga_kmr 65 59.72 ceb_fi 54 44.43 ha_uk 51 51.56 inh_mfe 48 24.07 hu_la 33 50.29 am_so 11 50.97 hr_hup 3 33.12
es_tk 64 40.68 bo_ceb 54 32.97 fa_ha 51 48.88 ca_mfe 48 38.2 hr_la 33 50.17 bn_ht 11 46.82 hup_nb 3 27.2
it_tk 64 41.45 bn_ceb 54 52.26 bg_ha 51 41.27 hy_mfe 48 46.13 la_ne 33 59.65 ha_pt 11 50.07 hi_hup 3 32.54
dz_pt 64 40.46 be_ceb 54 41.51 ha_ko 51 55.3 mfe_ne 48 53.64 la_nb 33 49.3 as_pt 10 75.72 he_hup 3 36.23
aeb_fa 64 68.26 ceb_et 54 42.89 es_ha 51 46.62 el_inh 47 30.03 hi_la 33 56.9 la_pt 10 45.23 hup_pt-br 3 30.27
fo_vi 63 48.79 ceb_eo 54 46.88 en_ha 51 57.12 fo_id 46 49.58 he_la 33 45.19 dz_pl 10 44.34 hup_my 3 46.37
nn_zh 63 60.29 az_ceb 54 44.36 el_ha 51 44.37 es_inh 46 28.37 la_pt-br 33 49.05 th_tk 10 42.32 hup_mt 3 31.76
bo_bs 63 43.99 am_ceb 54 45.32 ha_tr 51 44.45 inh_ru 46 25.83 la_mr 33 51.78 ro_tk 10 32.84 mt_nb 3 80.25
bs_km 63 62.47 ceb_ta 54 50.19 ar_ha 51 44.39 aeb_pt 45 76.27 af_ga 32 58.6 af_mfe 10 37.27 hi_mt 3 63.17
bo_km 63 47.71 af_ceb 54 43.24 ha_pl 51 47.55 af_gu 44 72.62 mfe_zh 31 39.96 la_zh 9 46.38 bi_ru 3 29.42
fr-ca_mfe 62 36.97 ast_ceb 54 37.96 ha_zh-tw 51 47.56 ckb_ha 44 46.68 nl_oc 30 33.26 pl_tk 9 34.85 lv_szl 3 55.01
ckb_tk 62 44.28 ceb_sl 54 44.1 ha_ja 51 55.31 ha_kmr 44 53.49 ro_tlh 30 24.69 sr_tk 9 37.66 ja_szl 3 39.65
tk_zh-cn 62 37.35 ceb_da 54 46.48 ha_sr 51 45.55 am_bs 43 65.81 af_lo 29 56.8 ht_ps 9 50.09 pt_szl 3 34.04
dz_el 62 45.12 ceb_oc 54 31.43 ha_sq 51 43.58 pt-br_szl 42 38.8 af_inh 29 31.05 gu_ug 9 66.01 arq_ht 2 40.03
fr_tk 61 38.95 ca_ceb 54 40.77 de_ha 51 43.72 mfe_sq 41 37.92 bs_ug 28 66.44 aeb_it 9 64.44 da_ht 2 41.83
dz_zh-tw 61 42.23 ceb_hy 54 50.01 ha_sk 51 47.48 af_nn 41 64.22 eu_ug 28 58.08 fr_inh 9 31.28 af_bi 2 32.58
dz_zh-cn 61 39.23 ceb_nb 54 46.76 ha_it 51 52.86 af_zh 40 59.43 kk_mg 27 47.26 fr-ca_pa 8 66.03 el_tk 2 45.76
fa_tk 60 44.95 ast_fi 54 56.77 ha_zh-cn 51 44.47 gl_ps 40 64.8 ceb_zh 27 44.84 be_ht 8 36.26 ht_mr 2 42.14
tk_zh-tw 60 41.8 fi_oc 54 36.88 ha_id 51 49.37 mfe_pt 39 44.45 ast_zh 27 44.92 af_ug 8 73.22 aeb_vi 2 55.35
ckb_ga 60 49.08 oc_uk 54 43 ha_ru 51 49.66 nn_te 39 57.31 oc_zh 27 32.03 ckb_sc 8 43.76 ckb_tt 1 88.13
gl_ig 59 40.04 fa_oc 54 41.87 ha_hy 51 51.5 bi_mn 39 44.58 gl_kn 27 66.78 ml_tl 8 51.02 as_ckb 1 47.01
id_tk 59 43.31 be_bo 54 37.13 ha_nl 51 54.74 af_arq 39 40.17 he_tlh 26 34.05 my_sc 8 54.14 fo_kk 1 43.46
aeb_de 59 68.22 am_bo 54 43.85 ha_hu 51 48.33 kn_ps 39 56.83 eu_si 25 68.28 am_tg 7 45.35 bi_pt 1 33.47
tk_tr 58 45.72 ast_bo 54 30.39 ha_hr 51 45.76 es_mfe 37 37.59 as_ro 25 64.36 ceb_pt 7 41.14 be_sh 1 56.24
hi_tk 58 47.52 bo_oc 54 23.06 ha_he 51 54.05 ast_pt 37 58.64 ml_ug 25 59 be_pa 7 81.32 am_ckb 1 78.39
my_tk 58 47.17 bn_oc 54 50 ha_pt-br 51 51.23 dz_sk 37 39.31 si_ug 25 58.91 pa_tl 7 65.73 ckb_tg 1 34.71
dz_fr 58 43.94 bg_oc 54 36.62 as_kmr 50 50.69 pt_ry 35 49.19 dz_fa 25 43.55 ps_tl 7 62.47 ceb_my 1 67.64
he_inh 58 29.27 am_be 54 57.44 el_oc 50 39.02 et_ps 35 57.86 eo_si 25 64.86 oc_pt 5 34.09 inh_vi 1 26.97
mfe_pt-br 57 43.86 ast_be 54 56.33 kk_nn 50 57.97 ml_si 34 59.55 en_tlh 24 39.46 bi_ro 5 26.19 eo_sh 1 50.2
ja_tk 57 42.41 be_oc 54 38.08 ga_mn 50 47.92 ko_mfe 33 42.88 el_tlh 24 34.64 hr_tk 5 38.05 is_sh 1 86.2
as_vi 57 61.09 ast_et 54 59.87 bi_vi 49 42.22 kk_mfe 33 46.4 sl_tlh 24 31.74 et_kn 5 58.4 nb_sh 1 90.78
as_cs 57 72.68 et_oc 54 36.08 bi_cs 49 31.51 af_tl 33 61.28 kmr_tlh 24 39.72 aeb_id 5 65.9
as_fr 57 67.12 es_oc 54 35 bi_lt 49 35.99 ga_gl 33 55.69 cs_inh 23 26.49 aeb_ja 5 63.36
Table 17: COMET-QE score and the count of sentences for all 4,765 language pairs in TED2025 (part VI).